### v1

# Hyperspherical Normalization for Scalable Deep Reinforcement Learning

Abstract
Scaling up the model size and computation has brought consistent performance improvements in supervised learning. However, this lesson often fails to apply to reinforcement learning (RL) because training the model on non-stationary data easily leads to overfitting and unstable optimization. In response, we introduce SimbaV2, a novel RL architecture designed to stabilize optimization by (i) constraining the growth of weight and feature norm by hyperspherical normalization; and (ii) using a distributional value estimation with reward scaling to maintain stable gradients under varying reward magnitudes. Using the soft actor-critic as a base algorithm, SimbaV2 scales up effectively with larger models and greater compute, achieving state-of-the-art performance on 57 continuous control tasks across 4 domains. The code is available at dojeon-ai.github.io/SimbaV2.
Machine Learning, ICML
[…]
Equal contribution
1 Introduction
Over the past decade, a scaling law has emerged as a cornerstone in deep learning, suggesting that increasing the model size, compute, and data consistently enhance performance (Kaplan et al., 2020; Dehghani et al., 2023). This paradigm has driven significant breakthroughs in supervised learning (SL), including large language models (Gemini et al., 2023; Achiam et al., 2023) and diffusion models (Ramesh et al., 2021; Rombach et al., 2022).
In contrast, scaling laws often fail to apply to reinforcement learning (RL) (Song et al., 2019; Li et al., 2023). Unlike SL where data distributions remain static, RL agents are trained under evolving data distributions and shifting objectives (Sutton & Barto, 2018). This inherent non-stationarity impedes scaling efforts, as increasing the model capacity or computational resources causes the model to easily overfit to earlier data and reduces the agent’s adaptability to new tasks (Lyle et al., 2022; Dohare et al., 2023).
Figure 1: Compute vs RL Performance. Soft Actor-Critic with SimbaV2 outperforms other RL algorithms, where performance scales as compute increases. The grey numbers below each dot indicate the update-to-data (UTD) ratio. SimbaV2, with UTD = 1, achieves a performance of 0.848, surpassing TD-MPC2 (0.749), the most computationally intensive version of Simba (0.818), and BRO (0.807). The results show normalized returns, averaged over $57$ continuous control tasks from MuJoCo, DMC, MyoSuite, and HumanoidBench, each trained on 1 million samples.
Figure 2: Benchmark Summary. (a) SimbaV2, with an update-to-data (UTD) ratio of 2, outperforms state-of-the-art RL algorithms across diverse continuous control benchmarks using fixed hyperparameters across all domains. (b) SimbaV2 delivers competitive performance in both online and offline RL while requiring significantly less training computation and offering faster inference times.
To address these challenges, the RL community has explored regularization techniques from SL. Methods such as weight decay (Farebrother et al., 2018) and dropout (Hiraoka et al., 2021) have been applied to reduce overfitting and improve generalization. Also, normalization techniques, such as layer normalization (Lei Ba et al., 2016; Lyle et al., 2023) and RL-specific variants (Bhatt et al., 2024; Lee et al., 2024c), have been employed to stabilize training by constraining the feature norm. More recently, periodic weight re-initialization (Nikishin et al., 2022; D’Oro et al., 2023; Schwarzer et al., 2023) has emerged as a strategy to address the instability in non-stationary optimization processes.
Despite their effectiveness, these approaches have notable limitations. Tuning regularizers is particularly challenging in RL’s dynamic settings; overly restrictive constraints impede training, while insufficient constraints fail to prevent overfitting (Henderson et al., 2018; Eimer et al., 2023). Moreover, normalization layers do not strictly regularize the growth of the parameter norm, resulting in varying effective learning rates (i.e., the gradient-to-parameter norm ratio) across layers and potentially destabilizing training (Lyle et al., 2024). Finally, periodic weight reinitialization becomes computationally prohibitive for large models or complex tasks (Sokar et al., 2023; Lee et al., 2024b).
In response, we present SimbaV2, a novel RL architecture designed to stabilize non-stationary optimization by constraining weight, feature, and gradient norms. Building on the Simba architecture (Lee et al., 2024c), which uses pre-layernorm residual blocks (Xiong et al., 2020) and weight decay (Krogh & Hertz, 1991) to control the weight and feature norm growth, SimbaV2 introduces two modifications:
  • Hyperspherical Normalization: We replace all layer normalization with hyperspherical normalization (i.e.,$\ell_{2}$-normalization) and project weights onto the unit-norm hypersphere after each gradient update. These changes ensure consistent effective learning rates across layers and eliminate the need for tuning weight regularization.
  • Distributional Value Estimation with Reward Scaling: To address unstable gradient norms from varying reward scales and outliers, we integrate a distributional critic (Bellemare et al., 2017) and apply reward scaling to ensure the unit variance of target values for both actor and critic.
Using soft actor-critic (Haarnoja et al., 2018) as a base, SimbaV2 stabilizes weight, feature, gradient norms, along with the effective learning rate throughout training (Section 5.2 and Figure 4). We evaluate SimbaV2 on four standard online RL benchmarks, MuJoCo (Todorov et al., 2012), DMC Suite (Tassa et al., 2018), MyoSuite (Caggiano et al., 2022), and HumanoidBench (Sferrazza et al., 2024), as well as D4RL MuJoCo benchmark (Fu et al., 2020) for offline RL. As indicated by results in Figures 1 and 2, SimbaV2 achieves state-of-the-art performance without algorithmic modifications or hyperparameter tuning. SimbaV2 also scales effectively with increased model size and computation, avoiding overfitting even in the absence of periodic reinitialization. We will release our code in the camera-ready version.
2 Related Work
2.1 Regularization in Deep Reinforcement Learning
Deep RL is particularly susceptible to overfitting due to its inherently non-stationary optimization process (Song et al., 2019). To address overfitting, researchers have adapted regularization techniques from SL, including weight decay (Farebrother et al., 2018), dropout (Hiraoka et al., 2021), and various normalization layers (Gogianu et al., 2021; Bjorck et al., 2021; Lyle et al., 2023; Gallici et al., 2024; Bhatt et al., 2024; Lee et al., 2024c; Elsayed et al., 2024; Palenicek et al., 2025). However, these methods often prove insufficient when scaling RL models, as larger computational resources and increased model sizes can easily exacerbate overfitting (Li et al., 2023; Nauman et al., 2024a).
To further scale computations and model sizes in RL, recent studies have explored periodic weight reinitialization strategies in RL to rejuvenate learning and escape local minima (D’Oro et al., 2023; Nauman et al., 2024b). These strategies include reinitializing weights to their initial distributions (Nikishin et al., 2022), interpolating between random and current weights (Xu et al., 2023; Schwarzer et al., 2023), utilizing momentum networks (Lee et al., 2024b), and selectively reinitializing dormant weights (Sokar et al., 2023). While promising, reinitialization has a notable limitation: it can lead to the loss of useful information and incur significant computational overhead as model size increases.
To address these limitations, we introduce SimbaV2, an architecture that explicitly constrains parameter, feature, and gradient norms throughout training. By constraining norms through hyperspherical normalization, SimbaV2 stabilizes an optimization process and eliminates the need for weight decay or periodic weight reinitialization.
2.2 Hyperspherical Representations in Deep Learning
Hyperspherical representations are widely used in deep learning across image classification (Salimans & Kingma, 2016; Liu et al., 2017b), face recognition (Wang et al., 2017; Liu et al., 2017a), variational autoencoders (Xu & Durrett, 2018), and contrastive learning (Chen et al., 2020). Using spherical embeddings is known to enhance feature separability (Wang & Isola, 2020), improving performance in tasks requiring precise discrimination. Recently, researchers have applied the hyperspherical normalization to intermediate features and weights to stabilize training in large-scale models such as diffusion models (Karras et al., 2024) and transformers (Loshchilov et al., 2024).
In this work, we apply hyperspherical normalization to RL. Unlike previous studies that focus on training the network on stationary data distributions with discrete inputs and outputs, we demonstrate their effectiveness on non-stationary data distributions with continuous inputs and outputs.
3 Preliminaries
As background, we briefly explain Soft Actor-Critic (SAC) algorithm (Haarnoja et al., 2018) and the Simba architecture (Lee et al., 2024c).
3.1 Soft Actor Critic
SAC is a prominent off-policy algorithm for continuous control. It aims to maximize both expected cumulative reward and policy entropy where $\tau=(o,a,r,o^{\prime})$ represents a transition tuple. SAC comprises a stochastic policy $\pi_{\theta}(a|o)$, a Q-function $Q_{\phi}(o,a)$, and an entropy coefficient $\alpha$ that balances reward and entropy.
The policy network is optimized to maximize the expected return while encouraging entropy which is formalized as:
$\mathcal{L}_{\pi}=\mathbb{E}_{\bar{a}\sim\pi_{\theta}}\left[\alpha\log\pi_{\theta}(\bar{a}|o)-Q_{\phi}(o,\bar{a})\right].$ (1)
The Q-function $Q_{\phi}(o,a)$ is trained to minimize the Bellman residual:
$\mathcal{L}_{Q}=(Q_{\phi}(o,a)-\left(r+\gamma Q_{\bar{\phi}}(o^{\prime},a^{\prime})-\alpha\log\pi_{\theta}(a^{\prime}|o^{\prime})\right))^{2},$ (2)
where $a^{\prime}\sim\pi_{\theta}(\cdot|o^{\prime})$, $\gamma\in[0,1)$ is the discount factor, and $Q_{\bar{\phi}}$ represents the target Q-network updated via an exponential moving average of $\phi$.
3.2 Simba Architecture
Simba (Lee et al., 2024c) is an RL architecture with normalization layers composed of the following stages:
Input Embedding. Given an input observation $\bm{o}_{t}\in\mathbb{R}^{|\mathcal{O}|}$, Simba applies Running Statistics Normalization (RSNorm) to normalize each dimension to zero mean and unit variance.
At each timestep $t$, the running mean $\bm{\mu}_{t}\in\mathbb{R}^{|\mathcal{O}|}$ and variance $\bm{\sigma}^{2}\in\mathbb{R}^{|\mathcal{O}|}$ are updated recursively as:
$\bm{\mu}_{t}=\bm{\mu}_{t-1}+\frac{1}{t}\bm{\delta}_{t},\quad\bm{\sigma}_{t}^{2}=\bm{\sigma}_{t-1}^{2}+\frac{1}{t}(\bm{\delta}_{t}^{2}-\bm{\sigma}_{t-1}^{2})$ (3)
where $\bm{\delta}_{t}=\bm{o}_{t}-\bm{\mu}_{t-1}$.
Given running statistics, the observation is normalized as:
$\bm{\bar{o}}_{t}=\text{RSNorm}(\bm{o}_{t})=\frac{\bm{o}_{t}-\bm{\mu}_{t}}{\sqrt{\bm{\sigma}_{t}^{2}+\epsilon}}.$ (4)
Then, the normalized observation, $\bar{\bm{o}}_{t}$, is embedded with a linear layer $\bm{W}_{h}^{0}\in\mathbb{R}^{|\mathcal{O}|\times d_{h}}$ defined as:
$\bm{h}_{t}^{0}=\bm{W}_{h}^{0}\bar{\bm{o}}_{t}.$ (5)
Latent Encoding. Next, the embedding $\bm{h}_{t}^{0}$ is encoded by a stack of $L$ residual blocks with pre-layer normalization. For $l\in\{1,\dots,L\}$, each of the $l$-th block is defined as:
$\bm{h}_{t}^{l}=\bm{h}_{t}^{l-1}+\text{MLP}(\text{LayerNorm}(\bm{h}_{t}^{l-1}))$ (6)
After the final block, the output is normalized again to obtain the latent feature:
$\bm{z}_{t}=\text{LayerNorm}(\bm{h}_{t}^{L}).$ (7)
Output Prediction: Finally, to predict the policy or Q-value, a linear layer $\bm{W}_{o}\in\mathbb{R}^{d_{h}\times d_{o}}$ maps $\bm{z}_{t}$ to:
$\bm{p}_{t}=\bm{W}_{o}\bm{z}_{t}.$ (8)
4 SimbaV2
SimbaV2 builds on Simba by adding constraints on weights, features, and gradients to enhance training stability, particularly when scaling to larger models and more computation. The modifications include:
  • LayerNorm $\bm{\rightarrow}$ $\ell_{2}$-Norm: Layer normalization is replaced with $\ell_{2}$-normalization, constraining intermediate features to have unit norm.
  • Linear $\bm{\rightarrow}$ Linear + Scaler: Standard linear layer is decoupled into a linear layer with weights constrained to a unit norm hypersphere, without a bias, and a learnable scaling vector that performs element-wise scaling.
  • Residual Connection $\bm{\rightarrow}$ LERP: Residual connection is replaced with a learnable linear interpolation (LERP), which combines raw and transformed features via a learnable interpolation vector.
  • MSE Loss $\bm{\rightarrow}$ KL-divergence Loss: MSE-based Bellman loss is replaced with KL-divergence loss, using a categorical critic (Bellemare et al., 2017).
  • No Reward Scaling $\bm{\rightarrow}$ Reward Scaling: Rewards are normalized with running statistics to stabilize the scale of both actor loss (Equation.1) and critic loss (Equation.2).
In the following subsections, we describe these modifications in detail.
Figure 3: SimbaV2 architecture. The input observation is first normalized using running statistics, then shifted along a new axis with a constant $c_{\text{shift}}$ to preserve magnitude information before being projected onto the unit hypersphere. The projected observation is passed through a linear layer, followed by a series of non-linear blocks and refined with LERP, serving as a residual connection. A final linear layer predicts the policy or value function.
4.1 Input Embedding
Following Simba, SimbaV2 first standardize the raw observations $\bm{o}_{t}\in\mathbb{R}^{|\mathcal{O}|}$ using RSNorm, yielding $\bar{\bm{o}}_{t}$. To further stabilize training, we map $\bar{\bm{o}}_{t}$ onto the unit hypersphere before applying a linear layer.
Shift + $\bm{\ell_{2}}$-Norm.
Direct $\ell_{2}$-normalization can discard magnitude information (e.g., $\bar{\bm{o}}_{t}=[1,0]$ and $\ [2,0]$ both map to $[1,0]$). To retain magnitude information, we embed $\bar{\bm{o}}_{t}$ into an ($|\mathcal{O}|+1$)-dimensional vector by concatenating a positive constant $c_{\text{shift}}>0$, then apply $\ell_{2}$-normalization:
$\widetilde{\bm{o}}_{t}\;=\;\ell_{2}\text{-Norm}(\bigl{[}\bar{\bm{o}}_{t};\,c_{\text{shift}}\bigr{]}).$ (9)
As illustrated in Figure 3, this additional coordinate encodes the original norm of $\bar{\bm{o}}_{t}$, preserving magnitude information.
Linear + Scaler. We then embed $\bm{\tilde{o}}_{t}$ using a linear layer $\bm{W}^{0}_{h}\in\mathbb{R}^{(|\mathcal{O}|+1)\times d_{h}}$ and a scaling vector $\boldsymbol{s}_{h}^{0}\in\mathbb{R}^{d_{h}}$ as:
$\bm{h}_{t}^{0}=\ell_{2}\text{-Norm}(\bm{s}_{h}^{0}\odot(\bm{W}_{h}^{0}\;\mathrm{Norm}(\bm{\tilde{o}}_{t})).$ (10)
where the $\ell_{2}$-normalization projects back to the hypersphere.
4.2 Feature Encoding
Starting from the initial hyperspherical embedding $\bm{h}_{t}^{0}$, we apply $L$ consecutive blocks of non-linear transformations. Each $l$-th block transforms $\bm{h}_{t}^{l}$ into $\bm{h}_{t}^{l+1}$ as follows:
MLP + $\bm{\ell_{2}}$-Norm. Each block uses an inverted bottleneck MLP (Vaswani, 2017) followed by $\ell_{2}$-normalization to project the output back onto the unit hypersphere.
$\bm{\tilde{h}}_{t}^{l}=\ell_{2}\text{-Norm}(\bm{W}_{h,2}^{l}\,\mathrm{ReLU}\bigl{(}(\bm{W}_{h,1}^{l}\,\bm{h}_{t}^{l})\odot\bm{s}^{l}_{h}\bigr{)}).$ (11)
where $\bm{W}_{h,1}^{l}\in\mathbb{R}^{4d_{h}\times d_{h}}$ and $\bm{W}_{h,2}^{l}\in\mathbb{R}^{d_{h}\times 4d_{h}}$ are weight matrices, and $\bm{s}^{l}_{h}\in\mathbb{R}^{4d_{h}}$ is a learnable scaling vector.
LERP + $\bm{\ell_{2}}$-Norm. We then linearly interpolate between the original input $\bm{h}_{t}^{l}$ and its non-linearly transformed output $\bm{\tilde{h}}_{t}^{l}$, followed by another $\ell_{2}$-normalization:
$\bm{h}_{t}^{l+1}=\ell_{2}\text{-Norm}((\bm{1}-\bm{\alpha}^{l})\odot\bm{h}_{t}^{l}+\bm{\alpha}^{l}\odot\bm{\tilde{h}}_{t}^{l}).$ (12)
where $\bm{1}\in\mathbb{R}^{d_{h}}$ and $\bm{\alpha}^{l}\in\mathbb{R}^{d_{h}}$ are one vector and a learnable interpolation vector, respectively.
LERP acts analogous to a learnable residual connection but can also be viewed as a first-order approximation of a Riemannian retraction on the hypersphere (Absil et al., 2008). Please refer to Appendix A.1 for further discussion.
4.3 Output Prediction
We use a linear layer to parameterize both the policy distribution and Q-value. Because Simba’s single Q-value estimate with an MSE-based Bellman loss is susceptible to outliers, we adopt a categorical critic with KL-divergence loss (Bellemare et al., 2017), which provides smoother gradients and more stable optimization (Imani & White, 2018).
Distributional Critic. We represent the Q-value as a categorical distribution over a discrete set of returns:
$\bigl{\{}\delta_{i}=G_{\min}+(i-1)\,\frac{G_{\max}-G_{\min}}{n_{\text{atom}}-1}\;\big{|}\;i=1,...,n_{\text{atom}}\bigr{\}},$ (13)
where $G_{\min}$ and $G_{\max}$ denote the minimum and maximum possible returns, and $n_{\text{atom}}$ is the number of discrete atoms.
Given the encoded representation $\bm{h}_{t}^{L}$, we compute unnormalized logits $\bm{z}_{t}\in\mathbb{R}^{|\mathcal{A}|\times n_{\text{atom}}}$ for all actions as follows:
$\bm{z}_{t}\;=\;\bm{W}_{o,2}(\bigl{(}\bm{W}_{o,1}\,\bm{h}_{t}^{L}\bigr{)}\,\odot\,\bm{s}_{o}),$ (14)
where $\bm{W}_{o,1}\in\mathbb{R}^{d_{h}\times d_{h}}$, $\bm{W}_{o,2}\in\mathbb{R}^{|\mathcal{A}|\times n_{\text{atom}}\times d_{h}}$, and $\bm{s}_{o}\in\mathbb{R}^{d_{h}}$ are trainable parameters.
For each action $\bm{a}\in\mathcal{A}$, the categorical probability is represented by applying the softmax function to $\bm{z}_{t,\bm{a}}\in\mathbb{R}^{n_{\text{atom}}}$:
$p_{t,\bm{a}}\;=\;\mathrm{softmax}\bigl{(}\bm{z}_{t,\bm{a}}\bigr{)}.$ (15)
The resulting Q-value is the expected return under $p_{t,\bm{a}}$:
$Q(\bm{o}_{t},\bm{a})=\sum_{i=1}^{n_{\text{atom}}}\delta_{i}\,p_{t,\bm{a},i}.$ (16)
Reward Bounding and Scaling. To use a categorical critic, we first bound the target returns within $[G_{\min},G_{\max}]$ and then scale the reward to maintain unit variance, ensuring stable gradients for both the actor and the critic. Unlike previous work (Schaul et al., 2021), which scaled the critic loss, we scale the reward itself, affecting both components simultaneously. Moreover, unlike observation normalization, we do not center the reward, as shifting the reward can alter the optimal policy in episodic tasks (Naik et al., 2024).
Given a reward $r_{t}$ at time $t$ and a discounted factor $\gamma$, we track a running discounted return:
$G_{t}\leftarrow(1-\gamma)G_{t-1}+r_{t}$ (17)
where $G_{t}$ is re-initialized to 00 at the start of each episode.
Then, we track the running variance of $G_{t}$, denoted as $\sigma^{2}_{t,G}$ and maintain a running maximum:
$G_{t,\max}\leftarrow\max(G_{t,\max},G_{t}).$ (18)
We then scale the reward as follows:
$\bar{r}_{t}\leftarrow\frac{r_{t}}{\max(\sqrt{\sigma_{t,G}^{2}+\epsilon},\;G_{t,\max}/G_{\max})}.$ (19)
This formula stabilizes gradients for both high-variance and low-variance returns, while thresholding with $G_{t,\max}/G_{\max}$ ensures target returns remain within $[G_{\min},G_{\max}]$.
4.4 Initialization and Update
In this subsection, we outline how weight matrix $\bm{W}$, scaler $\bm{s}$, and interpolation vector $\bm{\alpha}$ are initialized and updated.
Weight. All weight vectors are initialized orthogonally and then projected onto the unit hypersphere which forms an orthonormal basis. At each gradient step, we re-project them onto the unit sphere to maintain unit norm.
Formally, let $\bm{W}$ be the weight matrix before the update, and let $\mathcal{L}$ denote the loss function. The update rule is defined as:
$\bm{W}\leftarrow\ell_{2}\text{-Norm}(\bm{W}-\eta\frac{\partial\mathcal{L}}{\partial\bm{W}})$ (20)
[…]
Scaler. Following Loshchilov et al. (2024), we decouple the initialization scale of $\bm{s}$ from its learning dynamics by using two scalars, $\bm{s}_{\mathrm{init}}$ and $\bm{s}_{\mathrm{scale}}$. Although $\bm{s}$ is initialized to $\bm{s}_{\mathrm{scale}}$, it behaves as if it was initialized to $\bm{s}_{\mathrm{init}}$ during the forward pass by:
$\bm{s}\leftarrow\bm{s}_{\mathrm{scale}}\odot(\bm{s}_{\mathrm{init}}\oslash\bm{s}_{\mathrm{scale}})$ (21)
[…]
Interpolation vector. Analogous to the scaler, the interpolation vector, $\bm{\alpha}$, also has $\bm{\alpha}_{init}$ and $\bm{\alpha}_{scale}$. Following Loshchilov et al. (2024), we initialize $\bm{\alpha}_{init}=\mathbf{1}/(L+1)$ and $\bm{\alpha}_{scale}=\mathbf{1}/\sqrt{d_{h}}$, to preserve residual feature and gradually integrate non-linear features.
Figure 4: SimbaV2 vs. Simba Training Dynamics. We track 4 metrics during training to understand the learning dynamics of SimbaV2: (a) Average normalized return across tasks. (b) Weighted sum of $\ell_{2}$-norms of all intermediate features in critic. (c) Weighted sum of $\ell_{2}$-norms of all critic parameters (d) Weighted sum of $\ell_{2}$-norms of all gradients in critic (e) Effective learning rate (ELR) of the critic. On both environments, SimbaV2 maintains stable norms and ELR, while Simba exhibits divergent fluctuations.
5 Experiments
We now present a series of experiments designed to evaluate SimbaV2. Our investigation centers on three main setups:
  • Empirical Analysis (Section 5.2). Investigates whether SimbaV2 stabilizes the non-stationary optimization process and whether scaling model capacity and computation improves performance without overfitting.
  • Comparisons (Sections 5.3). Compare SimbaV2 against state-of-the-art RL algorithms.
  • Design Study (Section 5.4.) Conducts ablation studies on individual architectural component of SimbaV2.
5.1 Experimental Setup
Environment. A total of 57 continuous-control tasks are considered across 4 domains: MuJoCo (Todorov et al., 2012), DMC Suite (Tassa et al., 2018), MyoSuite (Caggiano et al., 2022), and HumanoidBench (Sferrazza et al., 2024). Also, two challenging subsets are defined for an empirical analysis: DMC-Hard (7 tasks involving dog and humanoid embodiments) and HBench-Hard (5 tasks: run, balance-simple, sit-hard, stair, walk).
Baselines. Comparisons include a broad range of deep RL algorithms, PPO (Schulman et al., 2017), SAC (Haarnoja et al., 2018), TD3 (Fujimoto et al., 2018) TD3+OFE (Ota et al., 2020), TQC (Kuznetsov et al., 2020), DreamerV3 (Hafner et al., 2023), TD7 (Fujimoto et al., 2023), TD-MPC2 (Hansen et al., 2023), Cross-Q (Bhatt et al., 2024), BRO (Nauman et al., 2024b), MAD-TD (Voelcker et al., 2024), MR.Q (Fujimoto et al., 2025), and Simba (Lee et al., 2024c). Whenever available, we report the results from the original paper; otherwise, we run the authors’ official code. In addition, to further compare performance before and after scaling, we evaluate BRO, Simba, and SimbaV2 under both low UTD ratios ($\leq 2$) and high UTD ratios ($\leq 8$). Additional details are described in Appendix E.
Metrics. To aggregate performance across diverse domains, each environment’s return is normalized to a near $[0,1)$ range. Specifically, MuJoCo performance is normalized by TD3 (Fujimoto et al., 2018); DMC returns are divided by 1000; MyoSuite scores use success rates; and HumanoidBench scores are normalized by their success score.
Training. If possible, we tried to closely follow Simba’s training configuration aiming to provide an apples-to-apples comparison. Unless otherwise specified, the actor and critic have hidden dimensions of 128 and 512, respectively (approximately 5M parameters). The model is trained for 1M environment steps, using a UTD ratio 2. We used an Adam (Kingma & Ba, 2014) optimizer without weight decay and set the batch size to 256. The learning rate is linearly decayed from $1\times 10^{-4}$ to $3\times 10^{-5}$. Full hyperparameter configurations are provided in Appendix C.
For the parameter scaling experiments, we focus on scaling the critic as prior studies suggest that scaling the actor has limited benefits (Nauman et al., 2024b; Lee et al., 2024c). Specifically, the critic’s hidden dimension is scaled from $\{128,256,512,1024\}$, growing is parameters from $0.3$M to $17.8$M. For computation scaling experiments, the UTD ratio is scaled from $\{1,2,4,8\}$ both with and without periodic weight reinitialization. We apply reinitialization every 500k update steps as suggested by Nauman et al. (2024b).
Table 1: Online RL. Average final performance after 1M environment steps, where $\pm$ captures a 95% confidence interval (CI) computed over all raw benchmark samples. For algorithms with only average scores for each task available, we approximate the CI using these averages ($\dagger$). Note that this estimation may be inaccurate. The highest performance is highlighted. Any performance that is not statistically worse than the highest performance (according to Welch’s $t$-test with significance level $0.05$) is highlighted.
Mujoco (5) DMC-Easy (21) DMC-Hard (7) MyoSuite (10) HBench (14) All (57)
Method TD3.Norm Return (1k) Return (1k) Success Rate Success.Norm -
(a) Low UTD ($\leq$ 2)
PPO (Schulman et al., 2017) 00.447 $\pm$ 0.2700 00.327 $\pm$ 0.1280 00.033 $\pm$ 0.0300 0- 0- 0-
SAC (Haarnoja et al., 2018) 00.934 $\pm$ 0.0910 00.762 $\pm$ 0.0940 00.136 $\pm$ 0.040 00.607 $\pm$ 0.0880 00.279 $\pm$ 0.0500 00.554 $\pm$ 0.0570
TD3 (Fujimoto et al., 2018) $0$$\pm$ 0.0000 0- 0- 0- 0- 0-
TD3+OFE (Ota et al., 2020) 01.322 $\pm$ 0.2630 0- 0- 0- 0- 0-
TQC (Kuznetsov et al., 2020) 01.137 $\pm$ 0.1250 0- 0- 0- 0- 0-
TD7 (Fujimoto et al., 2023) 01.570 $\pm$ 0.0300 00.689 $\pm$ 0.1340 00.182 $\pm$ 0.1370 00.356 $\pm$ 0.1260 00.289 $\pm$ 0.0830 00.617 $\pm$ 0.3580
TD-MPC2 (Hansen et al., 2023) 01.040 $\pm$ 0.115 0 00.889 $\pm$ 0.0640 00.465 $\pm$ 0.1390 00.650 $\pm$ 0.1480 00.710 $\pm$ 0.1490 00.749 $\pm$ 0.1680
CrossQ (Bhatt et al., 2024) 01.475 $\pm$ 0.1410 0- 0- 0- 0- 0-
MR.Q (Fujimoto et al., 2025) 01.448 $\pm$ 0.1560 00.868 $\pm$ 0.0260 00.723 $\pm$ 0.0610 0- 0- 0-
BRO (Nauman et al., 2024b) 01.101 $\pm$ 0.1820 00.861 $\pm$ 0.0360 00.693 $\pm$ 0.0660 00.714 $\pm$ 0.0760 00.468 $\pm$ 0.1070 00.731 $\pm$ 0.0390
Simba (Lee et al., 2024c) 01.147 $\pm$ 0.0770 00.864 $\pm$ 0.0240 00.706 $\pm$ 0.050 00.743 $\pm$ 0.0790 00.606 $\pm$ 0.0730 00.780 $\pm$ 0.0280
SimbaV2 (ours) 01.617 $\pm$ 0.1030 00.874 $\pm$ 0.0250 00.729 $\pm$ 0.0650 00.847 $\pm$ 0.0660 00.776 $\pm$ 0.0710 00.892 $\pm$ 0.0320
(b) High UTD ($\geq$ 8)
DreamerV3 (Hafner et al., 2023) 00.760 $\pm$ 0.0950 00.714 $\pm$ 0.1240 00.009 $\pm$ 0.0060 00.482 $\pm$ 0.1660 00.022 $\pm$ 0.0230 00.397 $\pm$ 0.2890
MAD-TD (Voelcker et al., 2024) 0- 0- 00.708 $\pm$ 0.0650 0- 0- 0-
BRO (Nauman et al., 2024b) 01.150 $\pm$ 0.2020 00.871 $\pm$ 0.0340 00.767 $\pm$ 0.0590 00.814 $\pm$ 0.0660 00.619 $\pm$ 0.1170 00.807 $\pm$ 0.0370
Simba (Lee et al., 2024c) 01.175 $\pm$ 0.1360 00.866 $\pm$ 0.0360 00.720 $\pm$ 0.0870 00.834 $\pm$ 0.0980 00.657 $\pm$ 0.0990 00.818 $\pm$ 0.0430
SimbaV2 (ours) 01.598 $\pm$ 0.1760 00.876 $\pm$ 0.0350 00.769 $\pm$ 0.0890 00.866 $\pm$ 0.0900 00.822 $\pm$ 0.0990 00.911 $\pm$ 0.0440
5.2 Empirical Analysis
Training Dynamics. To understand the training dynamics of SimbaV2, we measure the feature norm, weight norm, gradient norm, and the effective learning rate (ELR), defined as the ratio of the gradient norm to the weight norm (Kodryan et al., 2022; Lyle et al., 2024) (See Appendix G for details). We weighted average each metric across layers where weights correspond to each layer’s fraction of total parameters. Additionally, we divide the layers into encoder layers (all layers before the output prediction) and predictor layers (those after) to analyze their respective dynamics.
Figure 4 compares SimbaV2 and Simba on DMC-Hard and HBench-Hard. As shown in Figure 4.(b)-(d), Simba exhibits large, often divergent fluctuations in feature, weight, and gradient norms between the encoder and predictor. Consequently, Figure 4.(e) shows that the encoder’s ELR trending upward while the predictor’s ELR declines.
In contrast, SimbaV2 enforces tighter constraints, stabilizing norms and ELRs throughout training. Although certain parameters (e.g., scalers or interpolation vectors) can exceed the unit norm, the majority of parameters remain on the hypersphere, resulting in more robust optimization. A standalone visualization of SimbaV2 is in Appendix G.
Figure 5: Parameter Scaling. We scale the number of model parameters by increasing the width of the critic network. On DMC-Hard, both Simba and SimbaV2 benefit from increased model size. On HBench-Hard, however, Simba plateaus at larger model sizes, whereas SimbaV2 continues to improve.
Parameter Scaling. Next, we investigate whether SimbaV2’s stable training dynamics alleviate overfitting and enhance performance as the model size scales. In Figure 5, we increase the critic network’s parameters on DMC-Hard (left) and HBench-Hard (right). Both Simba and SimbaV2 gain from a larger model on DMC-Hard. However, on HBench-Hard, despite comparable performance at the smallest scale (0.3M), Simba plateaus at larger scales, whereas SimbaV2 continues to improve. This suggests that stabilizing training dynamics can effectively leverage a larger model.
Compute Scaling. Finally, we explore compute scaling, a major factor in improving sample efficiency in deep RL (Li et al., 2023). While increasing the UTD ratio can enhance sample efficiency, it also raises the risk of overfitting. Previous work has addressed this with ensembling (Chen et al., 2021b), periodic reinitialization (Lee et al., 2024a; D’Oro et al., 2023; Nauman et al., 2024b), or both (Kim et al., 2023). We investigate whether SimbaV2’s stable dynamics enable effective scaling without these additional methods.
Figure 6: Compute Scaling.. We scale compute by increasing the UTD ratio. We compare Simba and SimbaV2, both with and without periodic reset. Simba saturates at lower ratios without reset but improves with reset. In contrast, SimbaV2 scales smoothly even without reset, where using reset slightly degrades its performance.
Figure 6 shows the effect of varying the UTD ratio on DMC-Hard (left) and HBench-Hard (right), comparing Simba and SimbaV2 with and without reinitialization (solid lines: no reinitialization; dashed lines: reinitialization). In Simba, performance plateaus at a UTD ratio of 2 on DMC-Hard and 1 on HBench-Hard. When combined with reinitialization, but further improves with reinitialization, consistent with D’Oro et al. (2023). In contrast, SimbaV2 scales consistently as the UTD ratio increases, even without reinitialization. Notably, reinitialization slightly degrades SimbaV2’s performance, as it disrupts training and adds time to recover.
To assess the importance of hyperspherical normalization for UTD scaling, we test a variant in Appendix H that adds distributional critics and reward scaling but omits the normalization. This variant fails to scale at higher UTD ratios, demonstrating the role of hyperspherical normalization.
Table 2: Design Study. We report the final performance for each design choice in the online RL benchmarks, averaged over 3 random seeds. Performance changes relative to the default SimbaV2 are highlighted according to their percentile difference: mild positive changes $[0.02,0.05)$ and mild negative changes $(-0.05,-0.02]$ are highlighted lightly, damaging changes $(-0.1,-0.05]$ are highlighted moderately, and catastrophic changes $(-1.0,-0.1]$ are highlighted boldly.
Design Mujoco (5) DMC-Easy (21) DMC-Hard (7) MyoSuite (10) HBench (14) All (57)
(idx) Original $\rightarrow$ Changed TD3.Norm Return (1k) Return (1k) Success Rate Success.Norm -
Input Projection
(a) L2 Normalize $\rightarrow$ No L2 Normalize 01.370 $\pm$ 0.2200 00.779 $\pm$ 0.0530 $0$$\pm$ 0.0940 00.815 $\pm$ 0.1000 00.711 $\pm$ 0.1230 00.809 $\pm$ 0.0590
(b) Shifting $\rightarrow$ No Shifting 01.406 $\pm$ 0.2330 00.771 $\pm$ 0.0560 00.724 $\pm$ 0.0880 $0$$\pm$ 0.1050 $0$$\pm$ 0.1250 00.810 $\pm$ 0.0610
(c) $c_{shift}:3\rightarrow 1$ 01.558 $\pm$ 0.1670 00.862 $\pm$ 0.0390 00.718 $\pm$ 0.0890 00.870 $\pm$ 0.0850 00.791 $\pm$ 0.1300 00.888 $\pm$ 0.0580
(d) Shift Projection $\rightarrow$ Resize Projection 01.623 $\pm$ 0.1760 00.842 $\pm$ 0.0430 00.720 $\pm$ 0.0930 00.852 $\pm$ 0.0830 00.779 $\pm$ 0.1220 00.884 $\pm$ 0.0580
Output Prediction
(e) Categorical Loss $\rightarrow$ MSE Loss 01.343 $\pm$ 0.0970 00.868 $\pm$ 0.0340 00.708 $\pm$ 0.0970 00.757 $\pm$ 0.1300 00.767 $\pm$ 0.1290 00.846 $\pm$ 0.0620
(f) Reward Scaling $\rightarrow$ No Scaling 01.395 $\pm$ 0.1510 00.852 $\pm$ 0.0340 00.712 $\pm$ 0.0920 00.840 $\pm$ 0.0770 00.735 $\pm$ 0.0850 00.852 $\pm$ 0.0420
(g) Reward Bounding $\rightarrow$ No Bounding 01.620 $\pm$ 0.1420 00.824 $\pm$ 0.0330 00.733 $\pm$ 0.0720 00.805 $\pm$ 0.1210 00.787 $\pm$ 0.1280 00.868 $\pm$ 0.0510
(h) Soft Target $\rightarrow$ Hard Target 01.589 $\pm$ 0.1750 00.878 $\pm$ 0.0370 00.746 $\pm$ 0.0810 00.848 $\pm$ 0.0860 00.770 $\pm$ 0.0940 00.890 $\pm$ 0.0510
Initialization & Update
(i) LR Decay $\rightarrow$ No LR Decay 01.562 $\pm$ 0.1620 00.858 $\pm$ 0.0420 00.719 $\pm$ 0.0650 00.810 $\pm$ 0.1140 00.754 $\pm$ 0.1190 00.863 $\pm$ 0.0640
(j) $s_{init}:\sqrt{2}/\sqrt{d_{h}}\rightarrow 1$ 01.571 $\pm$ 0.1050 00.873 $\pm$ 0.0220 00.718 $\pm$ 0.0520 00.855 $\pm$ 0.0620 00.781 $\pm$ 0.0740 00.890 $\pm$ 0.0320
(k) $s_{scale}:\sqrt{2}/\sqrt{d_{h}}\rightarrow 1$ 01.594 $\pm$ 0.1020 00.870 $\pm$ 0.0250 00.706 $\pm$ 0.0550 00.836 $\pm$ 0.0530 00.789 $\pm$ 0.0720 00.887 $\pm$ 0.0330
(l) $\alpha_{init}:1/(L+1)\rightarrow 0.5$ 01.583 $\pm$ 0.1720 00.866 $\pm$ 0.0380 00.728 $\pm$ 0.0840 00.843 $\pm$ 0.0680 00.745 $\pm$ 0.1020 00.877 $\pm$ 0.0570
(m) $\alpha_{scale}:1/\sqrt{d_{h}}\rightarrow 1$ 01.520 $\pm$ 0.1770 00.856 $\pm$ 0.0340 00.714 $\pm$ 0.0890 00.875 $\pm$ 0.0790 00.792 $\pm$ 0.1250 00.885 $\pm$ 0.0590
SimbaV2 01.617 $\pm$ 0.1030 00.874 $\pm$ 0.0250 00.729 $\pm$ 0.0640 00.847 $\pm$ 0.0660 00.776 $\pm$ 0.0710 00.892 $\pm$ 0.0320
5.3 Online RL
Having observed SimbaV2’s scalability, we now compare it against standard model-free and model-based RL.
Table 1.(a) presents results at a UTD ratio below 2. SimbaV2 with UTD=2 attains an average normalized score of 0.892, exceeding the previous best of 0.780. Only except for DMC-Easy suite, SimbaV2 outperforms leading model-free (CrossQ (Bhatt et al., 2024), BRO (Nauman et al., 2024b), Simba (Lee et al., 2024c)) and model-based (TD-MPC2 (Hansen et al., 2023), Mr.Q (Fujimoto et al., 2025)) baselines, demonstrating superior sample efficiency.
Table 1.(b) evaluates higher UTD settings. Increasing SimbaV2’s UTD from 2 to 8 further elevates its average score from 0.892 to 0.911. SimbaV2 also surpasses BRO with UTD=10, which utilizes periodic reinitialization to avoid overfitting at high update rates. These consistent gains at larger UTD ratios underscore the efficacy of hyperspherical normalization in stabilizing training.
For offline RL, we simply add a behavioral cloning loss during training with using identical configurations to the online RL. Despite minimal changes, SimbaV2 performs competitively with existing baselines (Appendix D).
5.4 Design Study
Table 2 presents the results from ablation studies isolating the contributions of various architectural choices.
Input Projection. Projecting observations onto a hypersphere before passing them through the linear layer is crucial for performance (Table 2.(a)), where omitting this step leads to a significant performance drop. Equally important design is preserving the original magnitude during projection (Table 2.(b)). We also explore an alternative “resize” projection, where inputs are first divided by $c_{\rm{shift}}\sqrt{d_{h}}$ before being projected onto an $(n+1)$-dimensional hypersphere. The resize projection yields comparable performance as it can also retain magnitude information (Table 2.(d)).
Output Projection. Incorporating a distributional critic and reward scaling improves performance, especially in environments with high reward variance like MuJoCo (Table 2.(e)–(f)). Bounding target returns proves essential for easier tasks (Table 2.(g)), such as cartpole in the DMC-Easy suite (Table 25). Without bounding, consistent high returns can diminish return variance, and scaling returns push target values beyond the range of the distributional critic, leading to collapse in the TD loss.
Initialization & Update. Gradually decaying the learning rate is critical. Without decay, the model may struggle to refine its predictions during later training stages, as SimbaV2 maintains an effective constant learning rate throughout training (Table 2.(i)). Tuning initial scaler values has minimal impact on performance where the architecture remains stable by these changes (Table 2.(j)–(m)).
6 Lessons and Opportunities
Lessons. Historically, RL research has relied on complex regularizations to address overfitting and scalability issues (Klein et al., 2024). Our findings suggest that suitably chosen constraints, exemplified by SimbaV2, can simplify these design complexities while retaining strong performance.
Opportunities. Future opportunities include deploying SimbaV2 in real-world robotics (Hwangbo et al., 2019), where sample efficiency is crucial, and extending it to model-based (Hansen et al., 2023) or visual RL (Kostrikov et al., 2020). Furthermore, with increasing interest in RL for training large language models (Ouyang et al., 2022; Guo et al., 2025), the potential benefits of using stricter normalization for large models remain an exciting open question.
Acknowledgements
We would like to express our gratitude to Dongyoon Hwang and Hawon Jeong for their valuable feedback on this paper.
Appendix
Appendix A Architectural Details
A.1 LERP: A Retraction-based Approximation on Riemannian Manifolds.
During the feature encoding stage in SimbaV2, the input $\bm{h}$ and its non-linearly transformed output $\bm{\tilde{h}}$ are linearly interpolated using a learnable interpolation vector $\bm{\alpha}\in\mathbb{R}^{d_{h}}$:
$\bm{h}\leftarrow\ell_{2}\textrm{-Norm}((\bm{1}-\bm{\alpha})\odot\bm{h}+\bm{\alpha}\odot\bm{\tilde{h}}),$ (22)
followed by $\ell_{2}$-normalization.
Intuitively, this can be interpreted as a first-order (retraction-based) approximation of the Riemannian update formula on the hypersphere. This section provides a brief introduction to the differential geometry concepts that underpin the Riemannian optimization perspective of $\bm{\alpha}$. For brevity, we omit the mathematical definitions, derivations, and proofs here. The comprehensive introduction to differential geometry and Riemannian optimization can be found in Spivak (1970), Do Carmo & Flaherty Francis (1992), and Boumal (2023).
Let $\mathbb{S}_{n-1}$ denote the $n$-dimensional hypersphere embedded in $\mathbb{R}^{n}$, i.e., $\mathbb{S}_{n-1}=\{\bm{h}\in\mathbb{R}^{n}\mid\|\bm{h}\|_{2}=1\}$.
Manifold. A manifold $\mathcal{M}$ of dimension $n$ is a space that can locally be approximated by a Euclidean space $\mathbb{R}^{n}$. The simplest examples of a manifold include the open ball $U=\{\bm{x}\in\mathbb{R}^{n}\mid\|\bm{x}\|_{2}<r\}$ for $r\in\mathbb{R}_{>0}$, and the hypersphere $\mathbb{S}_{n-1}$ is also a manifold in $\mathbb{R}^{n}$.
Tangent Spaces. At each point $\bm{x}\in\mathcal{M}$, the tangent space $T_{\bm{x}}\mathcal{M}$ is an $n$-dimensional vector space that locally approximates $\mathcal{M}$ near $\bm{x}$. Tangent vectors generalize the concept of directional derivatives. For the hypersphere $\mathbb{S}_{n-1}$, the tangent space at a point $\bm{p}$ consists of all vectors orthogonal to $\bm{p}$:
$T_{\bm{p}}\mathbb{S}_{n-1}=\{\bm{h}\in\mathbb{R}^{n}\mid\langle\bm{p},\bm{h}\rangle=0\}$ (23)
[…]
Riemannian Metrics and Manifolds. The tangent space $T_{\bm{x}}\mathcal{M}$ is not inherently equipped with an inner product. A Riemannian metric $\rho$ provides a collection of inner products $\rho_{\bm{x}}(\cdot,\cdot):T_{\bm{x}}\mathcal{M}\times T_{\bm{x}}\mathcal{M}\to\mathbb{R}$ on the tangent spaces, $\rho:=(\rho_{\bm{x}})_{\bm{x}\in\mathcal{M}}$, which locally define the geometry of $\mathcal{M}$. A Riemannian manifold $(\mathcal{M},\rho)$ is a smooth manifold $\mathcal{M}$ equipped with such a metric. This enables us to define geometric notions such as distance, angle, length, volume, and curvature of manifold. For a detailed explanation of geometrics on Riemannian manifolds, refer to Lee (2006).
Exponential Mapping and Retraction. Under some conditions (Do Carmo & Flaherty Francis, 1992), the exponential map $\exp_{\bm{x}}:T_{\bm{x}}\mathcal{M}\to\mathcal{M}$ can be defined at a point $\bm{x}\in\mathcal{M}$. $\exp_{\bm{x}}(\bm{v})$ maps a tangent vector $\bm{v}\in T_{\bm{x}}\mathcal{M}$ to a point on the manifold along the geodesic from $\bm{x}$ in the direction of $\bm{v}$. Therefore, for small $t\in\mathbb{R}$, $\exp_{\bm{x}}(t\bm{v})$ represents the shortest path on $\mathcal{M}$ starting at $\bm{x}$ with initial direction $\bm{v}$. In Euclidean space $(\mathbb{R}^{n},\bm{I}_{n})$, the exponential map $\exp_{\bm{x}}(\bm{v})=\bm{x}+\bm{v}$ is simply defined as a straight path. In practice, for computational efficiency (e.g., the mappings do not have closed-form), we often approximate the exponential map $\exp_{\bm{x}}$ by a retraction (Absil et al., 2008) $R_{\bm{x}}$:
Definition A.1 (Retraction).
A retraction $R$ on a manifold $\mathcal{M}$ is a smooth map:
[…]
with the following properties:
$R_{\bm{x}}(\bm{0})=\bm{x}\qquad\text{and}\qquad(dR_{\bm{x}})_{\bm{0}}=\text{id}$
where $R_{\bm{x}}$ denotes the restriction of $R$ to $T_{\bm{x}}\mathcal{M}$, $(dR_{\bm{x}})_{\bm{0}}$ denotes the differential of $R_{\bm{x}}$ at $\bm{0}$, and id is the identity map.
Intuitively, a retraction $R_{\bm{x}}(\bm{v})$ provides a first-order approximation of the exponential map $\exp_{\bm{x}}(\bm{v})$ (Boumal, 2023). Figure 7 illustrates the difference between the exponential map and retraction on $\mathbb{S}_{2}$. For the hypersphere $\mathbb{S}_{n-1}$, the retraction of a tangent vector $\bm{\xi}\in T_{\bm{h}}\mathbb{S}_{n-1}$ onto $\mathbb{S}_{n-1}$ is given by (Absil et al., 2008):
$R_{\bm{h}}(\bm{\xi})=\ell_{2}\textrm{-Norm}(\bm{h}+\bm{\xi})=\frac{\bm{h}+\bm{\xi}}{\|\bm{h}+\bm{\xi}\|_{2}}$ (24)
Figure 7: Exponential Map vs. Retraction on 3-dimensional sphere. Comparison of the exponential map $\exp_{\bm{x}}$ and the retraction $R_{\bm{x}}$ on the $3$-dimensional sphere $\mathbb{S}_{2}$. The exponential mapping sends a tangent vector $\bm{g}\in T_{\bm{x}}\mathbb{S}_{2}$ exactly along the geodesic from $\bm{x}$ to a point on the manifold, while the retraction locally approximates this mapping to first order. Figure adapted from Sutti & Yueh (2024).
Riemannian Optimization. On Riemannian manifolds, gradient updates ideally follow the curved geodesics, rather than straight lines as in Euclidean space. To this end, Bonnabel (2013) introduce Riemannian SGD that generalizes SGD to Riemannian manifolds using exponential map:
$\bm{h}\leftarrow\exp_{\bm{h}}(-\alpha\bm{g})$ (25)
where $\alpha>0$ is the global learning rate and $\bm{g}\in T_{\bm{h}}\mathcal{M}$ denotes the Riemannian gradient.
In our case, $-(\bm{\tilde{h}}-\bm{h})$ can be viewed as the gradient $\bm{g}$ in the Euclidean space. Then, we project the gradient onto the tangent space $T_{\bm{h}}\mathbb{S}_{n-1}$:
$\bm{g}_{\mathrm{proj}}$ $=\bm{g}-\langle\bm{g},\bm{h}\rangle\bm{h}$ (26)
$=-(\bm{\tilde{h}}-\bm{h})-\langle-\bm{\tilde{h}}+\bm{h},\bm{h}\rangle\bm{h}$ (27)
$=-\bm{\tilde{h}}+\langle\bm{\tilde{h}},\bm{h}\rangle\bm{h}$ (28)
Applying the retraction $\exp_{\bm{h}}(-\alpha\bm{g}_{\text{proj}})\approx R_{\bm{h}}(-\alpha\bm{g}_{\text{proj}})$:
$\bm{h}$ $\leftarrow\ell_{2}\textrm{-Norm}\left(\bm{h}+\alpha(\bm{\tilde{h}}-\langle\bm{\tilde{h}},\bm{h}\rangle\bm{h})\right)$ (29)
$=\ell_{2}\textrm{-Norm}\left((1-\alpha\langle\bm{\tilde{h}},\bm{h}\rangle)\bm{h}+\alpha\bm{\tilde{h}}\right)$ (30)
Thus, the LERP operation in SimbaV2 can be interpreted as a retraction-based approximation of the Riemannian update rule on the hypersphere, where the learning rate $\alpha$ is replaced by a learnable vector $\bm{\alpha}$ and the inner product $\langle\bm{\tilde{h}},\bm{h}\rangle$ term is neglected. Also, Loshchilov et al. (2024) empirically show that neglecting the inner product term has no significant impact on performance.
A.2 Scaler Initialization
In our algorithm, the scaler $\bm{s}\in\mathbb{R}^{d_{h}}$ is a learnable vector that element-wise scales the output $\bm{z}$ of the linear layer:
$\bm{z}=\bm{s}\odot\bm{W}\bm{h}\in\mathbb{R}^{d_{h}}$ (31)
where $\bm{W}\in\mathbb{R}^{d_{h}\times n}$ is the weight matrix of the linear layer, and $\bm{h}\in\mathbb{R}^{n}$ is the input vector. To ensure that $\bm{z}$ (approximately) maintains unit norm at initialization, we initialize $\bm{s}$ as $\bm{s}=\sqrt{\frac{2}{d_{h}}}\cdot\bm{1}$. The following section provides the derivation for this initialization.
We assume that each normalized embedding $\bm{w}_{l}\in\mathbb{R}^{n}$ of $\bm{W}$, and a random $n$-dimensional normalized vector $\bm{h}\in\mathbb{R}^{n}$, are uniformly distributed on the $n$-dimensional hypersphere $\mathbb{S}_{n-1}$. Furthermore, we assume that the vectors $\bm{w}_{l}$ and $\bm{h}$ are mutually independent (Feller, 1991). We denote the angle between $\bm{w}_{l}$ and $\bm{h}$ by $\theta_{l}$, such that $\cos\theta_{l}=\bm{w}_{l}\cdot\bm{h}$ since $\|\bm{w}_{l}\|_{2}=\|\bm{h}\|_{2}=1$.
Distribution of the Cosine of the Angle. For simplicity, assume that $\bm{w}_{l}$ is fixed. Since $\bm{h}$ is uniformly distributed on the hypersphere, the distribution of the angle $\theta_{l}$ depends on the solid angle (Weisstein, 2005) subtended by $\bm{h}$ with respect to $\bm{w}_{l}$. The surface area $A_{n-2}$ of an $(n-1)$-dimensional hyperspherical cap (Li, 2010) leads to the probability density function $f(\theta_{l})$ (Cai et al., 2013):
$f(\theta_{l})=\frac{A_{n-1}}{S_{n-1}}=\frac{\frac{2\pi^{(n-1)/2}}{\Gamma(\frac{n-1}{2})}}{\frac{2\pi^{n/2}}{\Gamma(\frac{n}{2})}}\sin^{n-2}(\theta_{l})=\frac{\Gamma(\frac{n}{2})}{\sqrt{\pi}\Gamma(\frac{n-1}{2})}\sin^{n-2}(\theta_{l})$ (32)
where $\theta_{l}\in[0,\pi]$, $\Gamma$ is the gamma function and $S_{n-1}$ is the surface area of $\mathbb{S}_{n-1}=\frac{2\pi^{n/2}}{\Gamma(\frac{n}{2})}$ (Weisstein, 2002).
Norm of Output Vector. Let $\bm{z}=\bm{s}\odot\bm{W}\bm{h}\in\mathbb{R}^{d_{h}}$ be the output of the linear layer. Each element of $\bm{z}$ and $\bm{s}$, denoted by $\bm{z}_{l}$ and $\bm{s}_{l}$, respectively, corresponds to the scaled cosine of the angle $\theta_{l}$ between $\bm{w}_{l}$ and $\bm{h}$:
$\bm{z}=\bm{s}\odot\bm{W}\bm{h}=\begin{bmatrix}\bm{s}_{1}(\bm{w}_{1}\cdot\bm{h})\\ \bm{s}_{2}(\bm{w}_{2}\cdot\bm{h})\\ \vdots\\ \bm{s}_{d_{h}}(\bm{w}_{d_{h}}\cdot\bm{h})\end{bmatrix}=\begin{bmatrix}\bm{s}_{1}\cos\theta_{1}\\ \bm{s}_{2}\cos\theta_{2}\\ \vdots\\ \bm{s}_{d_{h}}\cos\theta_{d_{h}}\end{bmatrix}$ (33)
The expected squared norm of $\bm{z}$ is then given by:
$\mathbb{E}[\|\bm{z}\|_{2}^{2}]=\sum_{l=1}^{d_{h}}\bm{s}_{l}^{2}\mathbb{E}[\cos^{2}\theta_{l}]$ (34)
Using the trigonometric identity $\cos^{2}(\theta)=\frac{1+\cos(2\theta)}{2}$ and the following integrals:
[…]
where the second integral vanishes due to the symmetry of $\cos(2\theta)$ about $\theta=\frac{\pi}{2}$, we compute the expectation:
[…]
$=\frac{\Gamma(\frac{n}{2})}{2\sqrt{\pi}\Gamma(\frac{n-1}{2})}\times\frac{\sqrt{\pi}\Gamma(\frac{n-1}{2})}{\Gamma(\frac{n}{2})}=\frac{1}{2}$ (39)
Thus, by setting $\bm{s}_{l}=\sqrt{\frac{2}{d_{h}}}$ for all $\ell\in\{1,\cdots,d_{h}\}$, we expect that the expected norm $\mathbb{E}[\|\bm{z}\|_{2}^{2}]$ is $1$ at initialization.
Appendix B Implementation Details
Listings LABEL:implementation:scaler, LABEL:implementation:input_embedding and LABEL:implementation:mlp provide the Google JAX implementation of scaling vector (Section 4.4), input embedding (Section 4.1), and MLP block (Section 4.2), respectively.
1import flax.linen as nn
2
3class Scaler(nn.Module):
4 dim: int
5 init: float
6 scale: float
7
8 def setup(self):
9 self.scaler = self.param(
10 nn.initializers.constant(1.0 * self.scale),
11 self.dim,
12 )
13 self.forward_scaler = self.init / self.scale
14
15 def __call__(self, x: jnp.ndarray) -> jnp.ndarray:
16 return self.scaler * self.forward_scaler * x
Listing 1: A JAX implementation of Scaler (Section 4.4)
1import jax.numpy as jnp
2import flax.linen as nn
3
4class InputEmbedding(nn.Module):
5 observation_dim: int
6 hidden_dim: int
7 shift_const: float
8 input_scaler_init: float
9 input_scaler_scale: float
10
11 def setup(self):
12 self.obs_rms = RunningMeanStd(
13 shape=self.observation_dim
14 )
15 self.w0 = nn.Dense(
16 features=self.hidden_dim,
17 use_bias=False
18 )
19 self.input_scaler = Scaler(
20 dim=self.observation_dim,
21 init=input_scaler_init,
22 scale=input_scaler_scale
23 )
24
25 def __call__(self, observation: jnp.ndarray) -> jnp.ndarray:
26 # RSNorm
27 o = (observations - self.obs_rms.mean) / jnp.sqrt(
28 self.obs_rms.var + self.epsilon
29 )
30 # Shift + l2-Norm
31 new_axis = jnp.ones((o.shape[:-1] + (1,))) * self.shift_const
32 o = jnp.concatenate([o, new_axis], axis=-1)
33 o = l2normalize(o, axis=-1)
34 # Linear + Scaler
35 h = self.w0(o)
36 h = self.input_scaler(h)
37 h = l2normalize(h, axis=-1)
38 return h
Listing 2: A JAX implementation of Input Embedding (Section 4.1).
1import flax.linen as nn
2
3class SimbaV2Block(nn.Module):
4 hidden_dim: int
5 ffn_scaler_init: float
6 ffn_scaler_scale: float
7 alpha_scaler_init: float
8 alpha_scaler_scale: float
9
10 def setup(self):
11 self.w1 = nn.Dense(
12 features=4*self.hidden_dim,
13 use_bias=False
14 )(x)
15 self.mlp_scaler = Scaler(
16 dim=4*self.hidden_dim,
17 init=ffn_scaler_init,
18 scale=ffn_scaler_scale
19 )
20 self.w2 = nn.Dense(
21 features=self.hidden_dim,
22 use_bias=False
23 )
24 self.alpha = Scaler(
25 dim=self.hidden_dim,
26 init=alpha_scaler_init,
27 scale=alpha_scaler_scale
28 )
29
30 def __call__(self, x: jnp.ndarray) -> jnp.ndarray:
31 residual = x
32 # MLP + l2-Norm
33 x = self.w1(x)
34 x = self.mlp_scaler(x)
35 x = nn.relu(x)
36 x = self.w2(x)
37 x = l2normalize(x, axis=-1)
38 # LERP + l2-Norm
39 x = l2normalize(residual + self.alpha(x - residual), axis=-1)
40 return x
Listing 3: A JAX implementation of MLP block (Section 4.2).
Appendix C Hyperparameters
For all experiments, we use consistent hyperparameters across benchmarks. The default settings are listed in Table 3.
Table 3: Hyperparameters Table. The hyperparameters listed below are used consistently across all tasks using SimbaV2, unless stated otherwise. For the discount factor $\gamma$, we set it automatically using heuristics used by TD-MPC2 (Hansen et al., 2023).
Hyperparameter Notation Value
Input Shift constant $c_{\text{shift}}$ 3.0
Output Number of return bins $n_{\text{atoms}}$ $101$
Support of return $[G_{\min},G_{\max}]$ $[-5,5]$
Reward scaler epsilon $\epsilon$ $1\mathrm{e}{-8}$
Training Input scaler $(\boldsymbol{s}_{h,\text{init}}^{0},\boldsymbol{s}_{h,\text{scale}}^{0})$ $(\sqrt{2}/\sqrt{d_{h}},\sqrt{2}/\sqrt{d_{h}})$
MLP scaler $(\boldsymbol{s}_{h,\text{init}}^{l},\boldsymbol{s}_{h,\text{scale}}^{l})$ $(\sqrt{2}/\sqrt{4d_{h}},\sqrt{2}/\sqrt{4d_{h}})$
Output scaler $(\boldsymbol{s}_{o,\text{init}},\boldsymbol{s}_{o,\text{scale}})$ $(\sqrt{2}/\sqrt{d_{h}},\sqrt{2}/\sqrt{d_{h}})$
LERP vector $(\boldsymbol{\alpha}_{\text{init}},\boldsymbol{\alpha}_{\text{scale}})$ $(1/(L+1),1/\sqrt{d_{h}})$
Behavior cloning weight $\lambda$ Online: $0.0$
Offline: $0.1$
Common Discount factor $\gamma$ Heuristic (Hansen et al., 2023)
Replay buffer capacity - $1$M
Buffer sampling - Uniform
Batch size - $256$
Update-to-data (UTD) ratio - $2$
TD steps $k$ $1$
Actor Number of blocks $L$ 1
Hidden dimension $d_{h}$ 128
Initial temperature $\alpha_{0}$ $1\mathrm{e}{-2}$
Target entropy $\mathbb{H}^{*}$ $|\mathcal{A}|/2$
Critic Number of blocks $L$ 2
Hidden dimension $d_{h}$ 512
Number of atoms $n_{atoms}$ 101
Target critic momentum $\tau$ $5\mathrm{e}{-3}$
Clipped double Q - Has Failure Termination (Mujoco, HBench): True
No Failure Termination (DMC, MyoSuite): False
Optimizer Optimizer - Adam
Optimizer momentum $(\beta_{1},\beta_{2})$ (0.9, 0.999)
Weight Decay - 0.0
Learning rate init $\eta$ $1\mathrm{e}{-4}$
Learning rate final - $3\mathrm{e}{-4}$
Appendix D Offline RL
In this section, we assess whether the SimbaV2 architecture also provide benefits in offline RL, training from a stationary distribution. We adopt the minimalist offline RL method from (Fujimoto & Gu, 2021), where the behavioral cloning loss is integrated into the reinforcement learning objective. The objective is defined as:
$\pi\approx\arg\max_{\pi}\mathbb{E}_{(s,a)\sim D}\left[Q(s,\pi(s))-\lambda\left|\mathbb{E}_{s\sim D}[Q(s,\pi(s))]\right|\cdot(\pi(s)-a)^{2}\right]$ (40)
where we used $\lambda=0.1$, as in (Fujimoto et al., 2023), and no parameter tuning is performed.
D.1 Experimental Setup
Environment. We use 9 MuJoCo tasks from the D4RL (Fu et al., 2020) benchmark, covering 3 environments (HalfCheetah, Hopper, Walker2d) and 3 difficulty levels (Medium, Medium-Replay, Medium-Expert).
Baselines. We compare SimbaV2 against standard offline RL methods: Percentile BC, Decision Transformer (DT, (Chen et al., 2021a)), Diffusion Q-Learning (DQL, (Wang et al., 2022)), Implicit Diffusion Q-Learning (IDQL, (Hansen-Estruch et al., 2023)), Conservative Q-Learning (CQL, (Chen et al., 2021a)), TD3+BC (Fujimoto & Gu, 2021), Implicit Q-Learning (IQL, (Kostrikov et al., 2021)), Extreme Q-Learning ($\mathcal{X}$-QL, (Garg et al., 2023)), and TD7+BC (Fujimoto et al., 2023).
The results for Percentile BC, DT, DQL, and IDQL is from (Hansen-Estruch et al., 2023), while CQL, TD3+BC, IQL, $\mathcal{X}$-QL, and TD7 results come from (Fujimoto et al., 2023).
Metrics. Following the standard offline RL protocol (Fu et al., 2020), we normalize the score of each environment based on the expert trajectory in the dataset.
Training. We use the same training configuration as in online RL (Appendix C), with a learning rate decaying linearly from $1\times 10^{-4}$ to $1\times 10^{-5}$ over 100 epochs, and include an additional behavioral cloning loss.
D.2 Results
Table 4 reports the performance of SimbaV2 + BC, averaged over 10 random seeds.
Table 4: Offline RL. Average final performance on the D4RL mujoco benchmark, averaged over 10 trials. Methods are listed in chronological order. For a fair comparison, we used a unified hyperparameter configuration for each method. The highest performance is highlighted. Any performance that is not statistically worse than the highest performance (according to Welch’s $t$-test with significance level $0.05$) is highlighted. The environment’s average variance was used for the statistical test for methods without reported variance.
HalfCheetah Hopper Walker2d Average
Method m m-r m-e m m-r m-e m m-r m-e -
CQL (Kumar et al., 2020) 046.7$\pm$0.30 045.5$\pm$0.30 076.8$\pm$7.40 059.3$\pm$3.30 078.8$\pm$10.90 079.9$\pm$19.80 081.4$\pm$1.70 079.9$\pm$3.60 0108.5$\pm$1.20 073.0$\pm$2.70
Percent BC (Chen et al., 2021a) 048.4 040.6 092.9 056.9 075.9 0110.9 075.0 062.5 0109.0 074.0
DT (Chen et al., 2021a) 042.6 036.6 086.8 067.6 082.7 0110.9 074.0 066.6 0108.1 074.7
TD3+BC (Fujimoto & Gu, 2021) 048.1$\pm$0.10 044.6$\pm$0.40 093.7$\pm$0.90 059.1$\pm$3.00 052.0$\pm$10.60 098.1$\pm$10.70 084.3$\pm$0.80 081.0$\pm$3.40 0110.5$\pm$0.40 074.6$\pm$1.70
IQL (Kostrikov et al., 2021) 047.4$\pm$0.20 043.9$\pm$1.30 089.6$\pm$3.50 063.9$\pm$4.90 093.4$\pm$7.80 064.2$\pm$32.00 084.2$\pm$1.60 071.2$\pm$8.30 0108.9$\pm$1.40 074.1$\pm$3.80
DQL (Wang et al., 2022) 050.6 045.8 093.3 075.2 094.5 0102.1 083.4 086.7 0109.6 082.4
$\mathcal{X}$-QL (Garg et al., 2023) 047.4$\pm$0.10 044.2$\pm$0.70 090.2$\pm$2.70 067.7$\pm$3.60 082.0$\pm$14.90 092.0$\pm$10.00 079.2$\pm$4.00 061.8$\pm$7.70 0110.3$\pm$0.20 075.0$\pm$2.30
IDQL (Hansen-Estruch et al., 2023) 049.7 045.1 094.4 063.1 082.4 0105.3 080.2 079.8 0111.6 079.1
TD7+BC (Fujimoto et al., 2023) 058.0$\pm$0.40 053.8$\pm$0.80 0104.6$\pm$1.60 076.1$\pm$5.10 091.1$\pm$8.00 0108.2$\pm$4.80 091.1$\pm$7.80 089.7$\pm$4.70 0111.8$\pm$0.60 087.2$\pm$1.60
SimbaV2+BC (ours) 054.8$\pm$0.50 048.6$\pm$0.80 092.2$\pm$1.40 098.1$\pm$2.40 099.9$\pm$0.60 0106.2$\pm$1.50 082.7$\pm$10.30 087.7$\pm$2.10 0110.6$\pm$0.60 086.7$\pm$1.60
With minimal changes, SimbaV2 performs highly competitively with existing offline RL algorithms, with statistically significantly better performance on Hopper. Again, this experimental results reinforces the importance of architectural design over complex algorithmic modifications. We believe our architectural approach offers exciting future potential for bridging offline and online RL (Ball et al., 2023; Zhou et al., 2024).
Appendix E Baselines
PPO (Lillicrap, 2015). Proximal Policy Optimization (PPO) is an on-policy policy gradient method that constrains updated policies to remain proximal to the old policies to circumvent performance collapse. Results for Gym - MuJoCo and DMC were obtained from Fujimoto et al. (2025), which are averaged over $10$ seeds.
SAC (Haarnoja et al., 2018). Soft Actor-Critic (SAC) is an off-policy actor-critic algorithm in which the actor simultaneously maximizes expected return and entropy, encouraging both stability and exploration. Results for Gym-MuJoCo were obtained from Fujimoto et al. (2023). These scores are average over $10$ random seeds. For DMC, MyoSuite, and HBench tasks, we use the results from Lee et al. (2024c) which were obtained by running the official repository for 10 random seeds.
TD3 (Fujimoto et al., 2018). Twin Delayed DDPG (TD3) is an off-policy actor-critic algorithm that mitigates Q-overestimation bias via three key techniques: (i) clipped double Q-learning, (ii) delayed policy updates, (iii) target policy smoothing. Results for Gym-MuJoCo were obtained from Table 1 of Fujimoto et al. (2023). These scores are averaged over $10$ random seeds.
TD3+OFE (Ota et al., 2020). By replacing the encoder with an Online Feature Extractor (OFE)—trained via a dynamics prediction task to produce high-dimensional representations of observation-action pairs—TD3+OFE outperforms the original TD3 without requiring any hyperparameter adjustments. Results for Gym-MuJoCo were obtained from Table 1 of Fujimoto et al. (2023). These scores are averaged over $10$ random seeds. We attach these results into Table 1 by TD3-normalizing the scores as outlined in Appendix F.1.
TQC (Kuznetsov et al., 2020). Truncated Quantile Critic (TQC) proposes to truncate the return distribution of the distributional critics to flexibly balance between under- and overestimation bias of Q-value. Results for Gym-MuJoCo were taken directly from Table 1 of Fujimoto et al. (2023). We attach these results into Table 1 by TD3-normalizing the scores as described in Appendix F.1.
DreamerV3 (Hafner et al., 2023). DreamerV3 encodes sensory inputs into categorical representations to build a learned world model, enabling long-horizon behavior learning in its compact latent space. Results for Gym-MuJoCo and DMC were obtained from Fujimoto et al. (2025), which are averaged over $10$ seeds. For MyoSuite, and HBench tasks, we use the results from Lee et al. (2024c) which were obtained by running the official repository (https://github.com/SonyResearch/simba) over 3 random seeds.
TD7 (Fujimoto et al., 2023). TD7 improves TD3 by combining TD3 with four key improvements: (i) state-action representation learning (SALE), (ii) prioritized experience replay, (iii) policy checkpoints, and (iv) additional behavior cloning loss for offline RL. Results for Gym-MuJoCo and DMC were obtained from Fujimoto et al. (2025), which are averaged over $10$ seeds. For MyoSuite, and HBench tasks, we use the results from Lee et al. (2024c) which were obtained by running the official repository (https://github.com/SonyResearch/simba) over $5$ random seeds.
TD-MPC2 (Hansen et al., 2023). TD-MPC2 is a model-based algorithm that learns an implicit (decoder-free) world model through multiple dynamics prediction tasks and performs local trajectory optimization within the learned latent space. Results for Gym-MuJoCo and DMC were obtained from Fujimoto et al. (2025), which are averaged over $10$ seeds. For MyoSuite, and HBench tasks, we use the results from Lee et al. (2024c) which were obtained by running the official repository (https://github.com/SonyResearch/simba) over 3 random seeds.
CrossQ (Bhatt et al., 2024). CrossQ achieves superior performance and sample efficiency with low replay ratio, by removing target networks and employing careful batch normalization. Results for Gym-MuJoCo were obtained by running the official repository (https://github.com/adityab/CrossQ) for 10 random seeds.
iQRL (Scannell et al., 2024). Implicitly Quantized Reinforcement Learning (iQRL) is a representation learning technique of model-free RL that prevents representation collapse and improve sample-efficiency via latent quantization. For the DMC hard tasks, results averaged over 3 random seeds were obtained directly from the authors.
BRO (Nauman et al., 2024b). Bigger, Regularized, Optimistic (BRO) scales the critic network of SAC by integrating distributional Q-learning, optimistic exploration, and periodic resets. Results for Gym-MuJoCo and DMC Easy were obtained by running the official repository (https://github.com/naumix/BiggerRegularizedOptimistic) for 5 random seeds. For DMC hard, MyoSuite, and HBench tasks, we use the results from Lee et al. (2024c) which were obtained by running the official repository (https://github.com/SonyResearch/simba) over $5$ random seeds for HBench tasks and 10 random seeds for DMC hard and MyoSuite tasks. Unless stated otherwise, we set update-to-data (UTD) ratio to be $2$.
MAD-TD (Voelcker et al., 2024). Model-Augmented Data for Temporal Difference learning (MAD-TD) aims to stabilize high UTD training by mixing a small fraction $\alpha$ of model-generated on-policy data with real off-policy replay data. For the DMC hard tasks, results averaged over $10$ random seeds were obtained directly from the authors using the best algorithm setting ($\text{UTD}=8$, $\alpha=0.05$).
MR.Q (Fujimoto et al., 2025). Model-based Representations for Q-learning (MR.Q) is a model-free algorithm that uses model-based objectives, such as dynamics and reward prediction, to obtain rich representation for actor-critic agent. We use the results for Gym-MuJoCo and DMC from Fujimoto et al. (2025) which were obtained by running the official repository (https://github.com/facebookresearch/MRQ) over $10$ random seeds.
Simba (Lee et al., 2024c). SimBa is an architecture designed to scale up parameters in deep reinforcement learning by injecting a simplicity bias with observation normalizer, residual blocks, and layer normalizations. For Gym-MuJoCo, DMC, MyoSuite, and HBench tasks, we use the results from Lee et al. (2024c) which were obtained by running the official repository (https://github.com/SonyResearch/simba) over $15$ random seeds for DMC hard tasks and $10$ random seeds otherwise.
Table 5: Environment details. We list the episode length, action repeat for each domain, total environment steps, and performance metrics used for benchmarking SimbaV2.
Gym DMC MyoSuite HumanoidBench
Episode length $1,000$ $1,000$ $100$ $500$ $1,000$
Action repeat $1$ $2$ $2$ $2$
Effective length $1,000$ $500$ $50$ $250$ $500$
Total env. steps $1$M $1$M $1$M $1$M
Performance metric Average Return Average Return Average Success Average Return
Appendix F Environment Details
This section outlines the benchmark environments used in our evaluation. A complete list of all tasks from each benchmark, including their observation and action dimensions, is provided at the end of this section. Additionally, Table 5 outlines the episode length, action repeat, total number of environment steps, and performance metrics for each task domain.
F.1 Gym - MuJoCo
Gym (Brockman, 2016; Towers et al., 2024) is a suite of benchmark environments spanning finite MDPs to Multi-Joint dynamics with Contact (Todorov et al., 2012, MuJoCo) simulations. It offers a diverse range of tasks, including classic Atari games, small-scale tasks such as Toy Text and classic controls, as well as physics-based continuous robot control. For our experiments, we focus on 5 locomotion tasks within MuJoCo environments, which simulate complex physical interactions involving multi-body dynamics and contact forces. A complete list of these tasks is provided in Table 6. Note that we use the v4 version.
For comparison across different score scales of each task, all MuJoCo scores are normalized using TD3 and the random score for each task, as provided in TD7 (Fujimoto et al., 2023).
[…]
Task Random TD3
Ant-v4 $-70.288$ $3942$
HalfCheetah-v4 $-289.415$ $10574$
Hopper-v4 $18.791$ $3226$
Humanoid-v4 $120.423$ $5165$
Walker2d-v4 $2.791$ $3946$
F.2 DeepMind Control Suite
DeepMind Control Suite (Tassa et al., 2018, DMC) is a standard continuous control benchmarks, encompassing a variety of locomotion and manipulation tasks with varying levels of complexity. These tasks range from simple low-dimensional settings ($\mathcal{O}\in\mathbb{R}^{3}$, $\mathcal{A}\in\mathbb{R}^{1}$) to highly complex scenarios ($\mathcal{O}\in\mathbb{R}^{223}$, $\mathcal{A}\in\mathbb{R}^{38}$). Our evaluation includes 27 DMC tasks, divided into two categories: DMC-Easy&Medium and DMC-Hard. All Humanoid and Dog tasks are grouped as DMC-Hard, while the rest are fall under DMC-Easy&Medium. Comprehensive lists of DMC-Easy&Medium and DMC-Hard are available in Tables 7 and 8, respectively.
F.3 MyoSuite
MyoSuite (Caggiano et al., 2022) models human motor control using musculoskeletal simulations of the human elbow, wrist, and hand, focusing on physiologically accurate movements. It provides benchmarks for intricate real-world object manipulation, ranging from simple posing tasks to the simultaneous manipulation of two Baoding balls. Our evaluation focuses on 10 MyoSuite tasks involving the hand. As defined by the authors, each task is categorized as hard when the goal is randomized; otherwise the goal is fixed. The full list of MyoSuite tasks is presented in Table 9.
F.4 HumanoidBench
HumanoidBench (Sferrazza et al., 2024) serves as a high-dimensional simulated robot learning benchmark, leveraging the Unitree H1 humanoid robot equipped with dexterous hands. It encompasses a diverse set of whole-body control tasks, spanning from fundamental locomotion to complex human-like activities that require refined manipulation. In our experiments, we concentrate on 14 locomotion tasks. A comprehensive list of tasks is provided in Table 10.
Note that the locomotion tasks do not necessitate hand dexterity. Therefore, to reduce the complexity arising from high degrees of freedom (DoF) and complex dynamics, we streamline the environments setup by excluding the hands of humanoid. For example, in case of walk, this drastically declines the dimension of the observation and action spaces by approximately 66%.
walk Without hand With 2 hand
Observation dim $|\mathcal{O}|$ $51$ $151$
Action dim $|\mathcal{A}|$ $19$ $61$
DoF (body) $25$ $25$
DoF (two hands) 00 $50$
For comparison across different score scales of each task, all HumanoidBench scores are normalized using each task’s target success score provided by the authors and random score. Random scores are measured by the average undiscounted returns over $10$ episodes of random agent. Each measurement is repeated over $10$ seeds.
[…]
Task Random Target Success
h1-balance-simple $9.391$ $800$
h1-balance-hard $9.044$ $800$
h1-crawl $272.658$ $700$
h1-hurdle $2.214$ $700$
h1-maze $106.441$ $1200$
h1-pole $20.09$ $700$
h1-reach $260.302$ $12000$
h1-run $2.02$ $700$
h1-sit-simple $9.393$ $750$
h1-sit-hard $2.448$ $750$
h1-slide $3.191$ $700$
h1-stair $3.112$ $700$
h1-stand $10.545$ $800$
h1-walk $2.377$ $700$
Table 6: Gym-MuJoCo. We evaluate a total of $5$ continuous control tasks from the Gym-MuJoCo benchmark. Below, we provide a list of all the tasks considered. The baseline performance for each task is reported at $1$M environment steps.
Task Observation dim $|\mathcal{O}|$ Action dim $|\mathcal{A}|$
Ant-v4 $27$ $8$
HalfCheetah-v4 $17$ $6$
Hopper-v4 $11$ $3$
Humanoid-v4 $376$ $17$
Walker2d-v4 $17$ $6$
Table 7: DMC-Easy Complete List. We evaluate a total of $21$ continuous control tasks from the DMC-Easy benchmark. Below, we provide a list of all the tasks considered. The baseline performance for each task is reported at $1$M environment steps.
Task Observation dim $|\mathcal{O}|$ Action dim $|\mathcal{A}|$
acrobot-swingup $6$ $1$
ball-in-cup-catch $6$ $1$
cartpole-balance $5$ $1$
cartpole-balance-sparse $5$ $1$
cartpole-swingup $5$ $1$
cartpole-swingup-sparse $5$ $1$
cheetah-run $17$ $6$
finger-spin $9$ $2$
finger-turn-easy $12$ $2$
finger-turn-hard $12$ $2$
fish-swim $24$ $5$
hopper-hop $15$ $4$
hopper-stand $15$ $4$
pendulum-swingup $3$ $1$
quadruped-run $78$ $12$
quadruped-walk $78$ $12$
reacher-easy $6$ $2$
reacher-hard $6$ $2$
walker-run $24$ $6$
walker-stand $24$ $6$
walker-walk $24$ $6$
Table 8: DMC-Hard Complete List. We evaluate a total of $7$ continuous control tasks from the DMC-Hard benchmark. Below, we provide a list of all the tasks considered. The baseline performance for each task is reported at $1$M environment steps.
Task Observation dim $|\mathcal{O}|$ Action dim $|\mathcal{A}|$
dog-run $223$ $38$
dog-trot $223$ $38$
dog-stand $223$ $38$
dog-walk $223$ $38$
humanoid-run $67$ $24$
humanoid-stand $67$ $24$
humanoid-walk $67$ $24$
Table 9: MyoSuite Complete List. We evaluate a total of $10$ continuous control tasks from the MyoSuite benchmark including both fixed-goal and randomized-goal (hard) settings. Below, we provide a list of all the tasks considered. The baseline performance for each task is reported at $1$M environment steps.
Task Observation dim $|\mathcal{O}|$ Action dim $|\mathcal{A}|$
myo-key-turn $93$ $39$
myo-key-turn-hard $93$ $39$
myo-obj-hold $91$ $39$
myo-obj-hold-hard $91$ $39$
myo-pen-twirl $83$ $39$
myo-pen-twirl-hard $83$ $39$
myo-pose $108$ $39$
myo-pose-hard $108$ $39$
myo-reach $115$ $39$
myo-reach-hard $115$ $39$
Table 10: HumanoidBench Complete List. We evaluate a total of $14$ continuous control locomotion tasks from the HumanoidBench benchmark that simulates the UniTree H1 humanoid robot. Below, we provide a list of all the tasks considered. The baseline performance for each task is reported at $1$M environment steps.
Task Observation dim $|\mathcal{O}|$ Action dim $|\mathcal{A}|$
h1-balance-hard $77$ $19$
h1-balance-simple $64$ $19$
h1-crawl $51$ $19$
h1-hurdle $51$ $19$
h1-maze $51$ $19$
h1-pole $51$ $19$
h1-reach $57$ $19$
h1-run $51$ $19$
h1-sit-simple $51$ $19$
h1-sit-hard $64$ $19$
h1-slide $51$ $19$
h1-stair $51$ $19$
h1-stand $51$ $19$
h1-walk $51$ $19$
Appendix G Training Stability
In Section 5.2, we investigated the training dynamics of SimbaV2 on DMC-Hard and HBench-Hard via four metrics: feature norm, parameter norm, gradient norm, and effective learning rate (ELR) of neural networks. This section presents these standalone metrics for SimbaV2 to highlight its stable behavior throughout training.
Effective Learning Rate. We base our notion of ELR on the effective step size of Kodryan et al. (2022), omitting the global learning rate $\eta$ and using dimension-based weighting $w_{i}=\frac{|\bm{\theta}_{i}|}{\sum_{j=1}^{N}|\bm{\theta}_{j}|}$ instead of squared-parameter-norm weighting $w_{i}=\frac{\|\bm{\theta}_{i}\|^{2}}{\sum_{j=1}^{N}\|\bm{\theta}_{j}\|^{2}}$.
Definition G.1 (Effective Learning Rate).
Let $\bm{\theta}=\{\bm{\theta}_{i}\}_{i=1}^{N}$ be the parameter set of a neural network, and $\bm{g}_{i}$ be the back-propagated gradient associated with $\bm{\theta}_{i}$. The (total) effective learning rate $\mathrm{ELR}$ of the network is defined as:
$\mathrm{ELR}\triangleq\sqrt{\sum_{i=1}^{N}w_{i}\frac{\|\bm{g}_{i}\|^{2}}{\|\bm{\theta}_{i}\|^{2}}}$ (41)
where $w_{i}=\frac{|\bm{\theta}_{i}|}{\sum_{j=1}^{N}|\bm{\theta}_{j}|}$. Intuitively, our $\mathrm{ELR}$ measures the “effective” gradient step—per parameter dimension—before scaled by the global learning rate.
Metrics. To reflect dimensional contributions across layers, we also apply the same weighting $w_{i}$ when computing the feature norm, parameter norm, and gradient norm. For instance, our gradient norm is defined as:
$\|\bm{g}_{i}\|^{2}\triangleq\sum_{i=1}^{N}w_{i}\|\bm{g}_{i}\|_{2}^{2}$ (42)
where $\|\cdot\|_{2}$ is the standard $\ell_{2}$-norm (Frobenius norm $\|\cdot\|_{F}$ in case of matrices). Analogous expressions are applied for feature and parameter norms. We separate encoder layers (all layers preceding the output) from predictor layers (all layers after) to capture their distinct roles in the network. Average returns are normalized by maximum score $1000$ for DMC-Hard, by success and random scores for HBench-Hard (Appendix F.4).
Results. Figure 8 shows the tracked metrics over 1 million training steps. Certain features (e.g., logits) and parameters (e.g., scalers and interpolation vectors) may occasionally exceed unit norm, the overall parameter norms are tightly controlled (Figure 8.(b)-(c)), and gradient magnitudes are consistently balanced across modules (Figure 8.(d)). This leads to the consistent trend and scales of their ELRs over time. We hypothesize that this stable behavior contributes to underpins SimbaV2’s improved performance and scalability.
Figure 8: SimbaV2 Training Dynamics. We track 4 metrics during training to understand the learning dynamics of SimbaV2: (a) Average normalized return across tasks. (b) Weighted sum of $\ell_{2}$-norms of all intermediate features in critic. (c) Weighted sum of $\ell_{2}$-norms of all critic parameters (d) Weighted sum of $\ell_{2}$-norms of all gradients in critic (e) Effective learning rate (ELR) of the critic. On both environments, SimbaV2 maintains feature and parameter norms aligned, producing consistent gradient norms and ELRs.
Appendix H Scalability Effect of Hyperspherical Normalization
In Section 5.2, we observe that SimbaV2 consistently scales with an increasing update-to-data (UTD) ratio, even without reinitialization, while Simba saturates at a ratio of $2$. However, this raises the question of whether hyperspherical normalization is critical for UTD scaling. This section investigates the effectiveness of hyperspherical normalization in scalability.
Experimental Setup. In this experiment, we examine a “Simba-like” variant, named Simba+, which incorporates distributional critic and reward scaling but only excludes the hyperspherical normalization. In other words, Simba+ is identical to SimbaV2 except that it excludes hyperspherical normalization. On DMC-Hard tasks, we compare SimbaV2, Simba, and Simba+ under varying model sizes and UTD ratios to determine the role of hyperspherical normalization in scaling performance.
Result. Figure 9 shows the scaling results. In Figure 9 (left), all three methods benefit from increased model capacity, but Simba+ slightly underperforms at larger parameter counts. More critically, Simba and Simba+ both plateau when the UTD ratio surpasses $2$ in Figure 9 (right), while SimbaV2 continues to improve. These results confirm that hyperspherical normalization is truly indispensable for UTD scaling.
Figure 9: Performance Scaling under DMC-Hard. We compare SimbaV2, Simba+, and Simba as scaling the number of model parameters by increasing the critic network width and UTD ratio. Simba+ fails to scale effectively at higher UTD ratios, highlighting the essential role of the hyperspherical normalization for scalability.
Appendix I Complete UTD Scaling Results
I.1 Gym - MuJoCo
Table 11: Gym - MuJoCo UTD Scaling Results. Final average performance at 1M environment steps for each of the $5$ locomotion tasks in the Gym - MuJoCo benchmark. The number of evaluated random seeds for each update-to-data (UTD) ratio is $5$. The values in [brackets] represent a 95% bootstrap confidence interval. The aggregate mean, median and interquartile mean (IQM) are computed over the TD3-normalized score as described in Appendix F.1.
Task UTD = 1 UTD = 2 UTD = 4 UTD = 8
Ant-v4 7405 [7315, 7496] 7429 [7209, 7649] 7230 [6968, 7492] 6940 [6431, 7449]
HalfCheetah-v4 11425 [10798, 12052] 12022 [11640, 12404] 12007 [11458, 12557] 11592 [9956, 13229]
Hopper-v4 3579 [3311, 3847] 4053 [3928, 4178] 4003 [3647, 4359] 4151 [4033, 4269]
Humanoid-v4 7696 [4385, 11008] 10545 [10195, 10896] 11133 [10908, 11358] 11703 [11282, 12125]
Walker2d-v4 6069 [5724, 6414] 6938 [6691, 7185] 6804 [6459, 7148] 6163 [4522, 7804]
IQM 1.433 [1.225, 1.648] 1.637 [1.471, 1.788] 1.617 [1.402, 1.83] 1.581 [1.358, 1.82]
Median 1.468 [1.269, 1.625] 1.616 [1.491, 1.743] 1.615 [1.438, 1.809] 1.602 [1.377, 1.821]
Mean 1.418 [1.264, 1.569] 1.617 [1.513, 1.719] 1.62 [1.47, 1.773] 1.598 [1.419, 1.78]
Figure 10: Gym-MuJoCo UTD Scaling Learning Curves. Average episode return (1k) for the Gym-MuJoCo environment. Results are averaged over $5$ random seeds, and the shaded areas indicate 95% bootstrap confidence intervals.
I.2 Deepmind Control Suite - Easy
Table 12: DMC-Easy UTD Scaling Results. Final average performance at 1M environment steps for each of the $21$ tasks of the DMC-Easy benchmark. The number of evaluated random seeds for each update-to-data (UTD) ratio is provided $5$. The values in [brackets] represent a 95% bootstrap confidence interval. The aggregate mean, median and interquartile mean (IQM) are reported in units of 1k.
Task UTD = 1 UTD = 2 UTD = 4 UTD = 8
acrobot-swingup 413 [376, 450] 436 [391, 482] 458 [359, 558] 477 [438, 516]
ball-in-cup-catch 981 [977, 985] 982 [980, 984] 982 [979, 986] 982 [979, 985]
cartpole-balance 999 [999, 999] 999 [999, 999] 999 [999, 999] 999 [999, 999]
cartpole-balance-sparse 1000 [1000, 1000] 967 [904, 1030] 1000 [1000, 1000] 1000 [1000, 1000]
cartpole-swingup 881 [881, 881] 880 [876, 883] 880 [879, 881] 881 [880, 882]
cartpole-swingup-sparse 845 [843, 848] 848 [848, 849] 848 [848, 849] 841 [824, 858]
cheetah-run 917 [913, 920] 920 [918, 922] 902 [868, 937] 916 [912, 920]
finger-spin 940 [895, 985] 891 [810, 972] 762 [608, 915] 910 [790, 1030]
finger-turn-easy 951 [916, 987] 953 [925, 980] 954 [917, 992] 936 [857, 1014]
finger-turn-hard 928 [885, 972] 951 [925, 977] 902 [866, 939] 950 [910, 990]
fish-swim 818 [779, 856] 826 [806, 846] 815 [780, 850] 807 [778, 836]
hopper-hop 379 [224, 535] 290 [233, 348] 326 [243, 410] 317 [230, 404]
hopper-stand 845 [704, 986] 944 [926, 962] 781 [449, 1112] 932 [898, 967]
pendulum-swingup 817 [776, 858] 827 [805, 849] 820 [781, 859] 821 [784, 859]
quadruped-run 931 [922, 940] 935 [928, 943] 943 [936, 949] 935 [930, 940]
quadruped-walk 962 [955, 970] 962 [955, 969] 964 [958, 971] 965 [958, 972]
reacher-easy 963 [927, 1000] 983 [979, 986] 975 [958, 992] 983 [981, 985]
reacher-hard 975 [971, 980] 967 [946, 987] 976 [972, 980] 974 [970, 978]
walker-run 813 [806, 819] 817 [812, 821] 821 [819, 823] 802 [774, 831]
walker-stand 986 [980, 992] 987 [984, 990] 988 [984, 992] 987 [984, 991]
walker-walk 977 [976, 979] 976 [974, 978] 976 [973, 979] 976 [972, 981]
IQM 0.928 [0.906, 0.948] 0.933 [0.918, 0.948] 0.925 [0.9, 0.946] 0.935 [0.91, 0.956]
Median 0.876 [0.834, 0.912] 0.875 [0.846, 0.904] 0.866 [0.818, 0.905] 0.878 [0.835, 0.917]
Mean 0.873 [0.838, 0.905] 0.874 [0.848, 0.897] 0.861 [0.823, 0.896] 0.876 [0.841, 0.908]
Figure 11: DMC-Easy UTD Scaling Learning Curves. Average episode return for the DMC-Easy environment. Results are averaged over $5$ random seeds, and the shaded areas indicate 95% bootstrap confidence intervals.
I.3 Deepmind Control Suite - Hard
Table 13: DMC-Hard UTD Scaling Results. Final average performance at 1M environment steps for each of the $7$ tasks of the DMC-Hard benchmark. The number of evaluated random seeds for each update-to-data (UTD) ratio is provided $5$. The values in [brackets] represent a 95% bootstrap confidence interval. The aggregate mean, median and interquartile mean (IQM) are reported in units of 1k.
Task UTD = 1 UTD = 2 UTD = 4 UTD = 8
dog-run 477 [429, 525] 562 [516, 608] 655 [620, 691] 555 [523, 587]
dog-stand 967 [959, 974] 981 [977, 985] 967 [960, 974] 972 [967, 976]
dog-trot 850 [810, 890] 861 [772, 950] 846 [782, 910] 898 [888, 909]
dog-walk 921 [912, 930] 935 [927, 944] 923 [905, 941] 949 [945, 953]
humanoid-run 183 [164, 203] 194 [182, 207] 272 [230, 313] 253 [228, 278]
humanoid-stand 660 [585, 734] 916 [886, 945] 928 [926, 930] 933 [924, 941]
humanoid-walk 568 [533, 603] 651 [590, 713] 818 [751, 885] 819 [762, 877]
IQM 0.713 [0.598, 0.809] 0.808 [0.725, 0.88] 0.851 [0.755, 0.916] 0.849 [0.727, 0.924]
Median 0.666 [0.563, 0.774] 0.729 [0.655, 0.81] 0.771 [0.678, 0.868] 0.767 [0.652, 0.861]
Mean 0.669 [0.581, 0.753] 0.729 [0.663, 0.791] 0.769 [0.687, 0.845] 0.759 [0.67, 0.84]
Figure 12: DMC-Hard UTD Scaling Learning Curves. Average episode return for the DMC-Hard environment. Results are averaged over $5$ random seeds, and the shaded areas indicate 95% bootstrap confidence intervals.
I.4 MyoSuite
Table 14: MyoSuite UTD Scaling Results. Final average performance at 1M environment steps across each of the $10$ continuous control tasks in the MyoSuite benchmark, including both fixed-goal and randomized-goal (hard) settings. The number of evaluated random seeds for each update-to-data (UTD) ratio is $5$. The values in [brackets] represent a 95% bootstrap confidence interval. Performance is measured by the average success rate of each task.
Task UTD = 1 UTD = 2 UTD = 4 UTD = 8
myo-pen-twirl-hard 76.0 [57.8, 94.2] 93.0 [88.8, 97.2] 92.0 [84.7, 99.3] 98.0 [94.1, 101.9]
myo-pen-twirl 100.0 [100.0, 100.0] 100.0 [100.0, 100.0] 100.0 [100.0, 100.0] 100.0 [100.0, 100.0]
myo-key-turn-hard 46.0 [8.5, 83.5] 62.0 [42.7, 81.3] 70.0 [34.9, 105.1] 80.0 [49.6, 110.4]
myo-key-turn 80.0 [40.8, 119.2] 100.0 [100.0, 100.0] 100.0 [100.0, 100.0] 100.0 [100.0, 100.0]
myo-obj-hold-hard 100.0 [100.0, 100.0] 98.0 [95.4, 100.6] 98.0 [94.1, 101.9] 92.0 [84.7, 99.3]
myo-obj-hold 100.0 [100.0, 100.0] 100.0 [100.0, 100.0] 100.0 [100.0, 100.0] 100.0 [100.0, 100.0]
myo-pose-hard 0.0 [0.0, 0.0] 0.0 [0.0, 0.0] 0.0 [0.0, 0.0] 0.0 [0.0, 0.0]
myo-pose 100.0 [100.0, 100.0] 100.0 [100.0, 100.0] 100.0 [100.0, 100.0] 100.0 [100.0, 100.0]
myo-reach-hard 92.0 [84.7, 99.3] 94.0 [87.3, 100.7] 98.0 [94.1, 101.9] 96.0 [91.2, 100.8]
myo-reach 100.0 [100.0, 100.0] 100.0 [100.0, 100.0] 100.0 [100.0, 100.0] 100.0 [100.0, 100.0]
IQM 96.9 [85.8, 100.0] 99.0 [96.8, 100.0] 99.2 [96.2, 100.0] 99.6 [96.9, 100.0]
Median 79.0 [67.0, 91.0] 84.5 [78.0, 93.0] 87.0 [76.0, 98.0] 88.0 [77.0, 98.0]
Mean 79.4 [68.6, 88.8] 84.7 [78.3, 90.6] 85.8 [76.4, 94.0] 86.6 [77.4, 94.6]
Figure 13: MyoSuite UTD Scaling Learning Curves. Average episode success rate (%) for the MyoSuite environment. Results are averaged over $5$ random seeds, and the shaded areas indicate 95% bootstrap confidence intervals.
I.5 Humanoid Bench
Table 15: HumanoidBench UTD Scaling Results. Final average performance at 1M environment steps for each of the $14$ locomotion tasks in the HumanoidBench benchmark. The number of evaluated random seeds for each update-to-data (UTD) ratio is $5$. The values in [brackets] represent a 95% bootstrap confidence interval. The aggregate mean, median and interquartile mean (IQM) are computed over the success normalized score as described in Appendix F.4.
Task UTD = 1 UTD = 2 UTD = 4 UTD = 8
h1-sit-hard-v0 681 [506, 857] 679 [548, 811] 719 [664, 773] 810 [784, 836]
h1-walk-v0 732 [522, 941] 845 [840, 850] 846 [841, 851] 844 [840, 847]
h1-stair-v0 473 [444, 503] 493 [467, 518] 546 [541, 550] 532 [512, 552]
h1-run-v0 247 [152, 342] 415 [307, 524] 318 [176, 461] 425 [293, 558]
h1-balance-simple-v0 806 [773, 839] 723 [651, 795] 775 [719, 831] 813 [797, 828]
h1-pole-v0 769 [758, 780] 791 [785, 797] 799 [780, 817] 827 [787, 868]
h1-slide-v0 412 [279, 544] 487 [404, 571] 544 [500, 588] 534 [505, 563]
h1-balance-hard-v0 135 [111, 160] 143 [128, 157] 128 [118, 139] 167 [157, 178]
h1-sit-simple-v0 873 [868, 879] 875 [870, 880] 908 [861, 955] 867 [839, 894]
h1-maze-v0 350 [332, 368] 313 [287, 340] 343 [327, 359] 338 [325, 351]
h1-crawl-v0 923 [884, 962] 946 [933, 959] 939 [927, 951] 954 [923, 984]
h1-hurdle-v0 193 [171, 215] 202 [167, 236] 244 [230, 259] 246 [206, 287]
h1-reach-v0 4166 [3706, 4627] 3850 [3272, 4427] 4003 [3614, 4392] 3449 [2541, 4358]
h1-stand-v0 771 [669, 873] 814 [770, 857] 765 [695, 835] 855 [828, 882]
IQM 0.734 [0.574, 0.887] 0.799 [0.685, 0.905] 0.813 [0.657, 0.958] 0.873 [0.706, 1.002]
Median 0.71 [0.612, 0.859] 0.781 [0.691, 0.863] 0.782 [0.665, 0.914] 0.824 [0.699, 0.944]
Mean 0.737 [0.637, 0.836] 0.776 [0.704, 0.846] 0.791 [0.687, 0.893] 0.822 [0.72, 0.92]
Figure 14: Humanoidbench UTD Scaling Learning Curves. Average episode return for the HumanoidBench environment. Results are averaged over $5$ random seeds, and the shaded areas indicate 95% bootstrap confidence intervals. The black dotted line indicates the success score of each tasks (Appendix F.4)
Appendix J Complete Main Results
This section provides learning curves and final performance for each online RL task across the evaluated algorithms.
Learning Curve. For visibility of learning curve, we focus on DreamerV3 (Hafner et al., 2023), TD7 (Fujimoto et al., 2023), TD-MPC2 (Hansen et al., 2023), MR.Q (Fujimoto et al., 2025), and Simba (Lee et al., 2024c) as main baselines, selected for their strong performance and community adoption. We omit curves for algorithms with unavailable raw samples at each task.
Confidence Interval. The light-colored area in the figures and the gray-shaded, bracketed terms in the tables represent 95% bootstrap confidence intervals. For each task evaluated over $n$ random seeds, the $95\%$ bootstrap confidence interval $\mathrm{CI}$ is computed as:
$\mathrm{CI}=\left[\mu-1.96\times\frac{\sigma}{\sqrt{n}},\mu+1.96\times\frac{\sigma}{\sqrt{n}}\right]$
where $\mu$ and $\sigma$ are the sample mean and standard deviation (with Bessel’s correction) of the evaluation, respectively. For aggregated scores (mean, median, and interquartile mean), confidence intervals are computed over all $n\times T$ raw samples, where $n$ and $T$ are the number of evaluated random seeds and tasks in the benchmark, respectively. For algorithms with only average scores for each task available, we approximate the CI of aggregated scores using these averages (denoted with gray-colored $\dagger$). We caution that this estimation may be inaccurate.
J.1 Gym - MuJoCo
Table 16: Gym - MuJoCo. Final average performance at 1M environment steps for each of the $5$ locomotion tasks in the Gym - MuJoCo benchmark. The number of evaluated random seeds for each algorithm is provided in Appendix E. The values in [brackets] represent a 95% bootstrap confidence interval. The aggregate mean, median and interquartile mean (IQM) are computed over the TD3-normalized score as described in Appendix F.1.
Task DreamerV3 TD7 TD-MPC2 MR.Q Simba SimbaV2
Ant-v4 1947 [1076, 2813] 8509 [8168, 8844] 4751 [2988, 6145] 6989 [6203, 7617] 5882 [5354, 6411] 7429 [7209, 7649]
HalfCheetah-v4 5502 [3717, 7123] 17433 [17301, 17559] 15078 [14065, 15932] 13305 [11841, 14140] 9422 [8745, 10100] 12022 [11640, 12404]
Hopper-v4 2666 [2106, 3210] 3511 [3236, 3736] 2081 [1197, 2921] 2684 [2154, 3269] 3231 [3004, 3458] 4054 [3929, 4179]
Humanoid-v4 4217 [2785, 5523] 7428 [7304, 7553] 6071 [5770, 6333] 7259 [5080, 9336] 6513 [5634, 7392] 10546 [10195, 10897]
Walker2d-v4 4519 [3692, 5244] 6096 [5621, 6547] 3008 [1706, 4321] 6629 [5816, 7493] 4290 [3864, 4716] 6938 [6691, 7185]
IQM 0.720 [0.620, 0.850] 1.540 [1.500, 1.580] 1.050 [0.890, 1.190] 1.450 [1.270, 1.580] 1.114 [1.043, 1.200] 1.637 [1.470, 1.791]
Median 0.810 [0.580, 0.930] 1.550 [1.450, 1.630] 1.180 [0.830, 1.220] 1.420 [1.190, 1.710] 1.143 [1.063, 1.227] 1.616 [1.49, 1.744]
Mean 0.760 [0.670, 0.860] 1.570 [1.540, 1.600] 1.040 [0.920, 1.150] 1.390 [1.270, 1.490] 1.147 [1.075, 1.223] 1.617 [1.513, 1.718]
Task PPO SAC TD3 TD3+OFE TQC CrossQ BRO
Ant-v4 1584 [1360, 1815] 3681 [3175, 4187] 3942 [2912, 4972] 7398 [7280, 7516] 3582 [2489, 4675] 6980 [6834, 7126] 7027 [6710, 7343]
HalfCheetah-v4 1744 [1523, 2118] 10484 [9825, 11143] 10574 [9677, 11471] 13758 [13214, 14302] 12349 [11471, 13227] 12893 [11771, 14015] 13747 [12621, 14873]
Hopper-v4 3022 [2633, 3339] 2785 [2151, 3419] 3226 [2911, 3541] 3121 [2615, 3627] 3526 [3302, 3750] 2467 [1855, 3079] 2122 [1655, 2588]
Humanoid-v4 477 [436, 518] 4090 [4545, 5273] 5165 [5020, 5310] 6032 [5698, 6366] 6029 [5498, 6560] 10480 [10307, 10653] 4757 [3139, 6376]
Walker2d-v4 2487 [1907, 3022] 4314 [4058, 4570] 3946 [3654, 4238] 5195 [4683, 5707] 5321 [4999, 5643] 6257 [5277, 7237] 3432 [2064, 4801]
IQM 0.410 [0.110, 0.834] 0.930 [0.812, 1.059] 1.000 [1.000, 1.000] 1.261 [1.035, 1.680] 1.143 [0.971, 1.290] 1.565 [1.394, 1.710] 1.071 [0.828, 1.333]
Median 0.412 [0.071, 0.936] 0.935 [0.787, 1.093] 1.000 [1.000, 1.000] 1.293 [0.967, 1.861] 1.163 [0.910, 1.349] 1.489 [1.317, 1.643] 1.071 [0.884, 1.322]
Mean 0.447 [0.186, 0.725] 0.934 [0.846, 1.027] 1.000 [1.000, 1.000] 1.322 [1.090, 1.615] 1.137 [1.012, 1.261] 1.475 [1.330, 1.608] 1.101 [0.927, 1.278]
Figure 15: Gym-MuJoCo Learning Curves. Average episode return (1k) for the Gym-MuJoCo environment. Results are averaged over random seeds of each algorithm, and the shaded areas indicate 95% bootstrap confidence intervals.
J.2 Deepmind Control Suite - Easy
Table 17: DMC Easy. Final average performance at 1M environment steps for each of the $21$ tasks of the DMC Easy benchmark. The number of evaluated random seeds for each algorithm is provided in Appendix E. The values in [brackets] represent a 95% bootstrap confidence interval. The aggregate mean, median and interquartile mean (IQM) are reported in units of 1k.
Task DreamerV3 TD7 TD-MPC2 MR.Q BRO Simba SimbaV2
acrobot-swingup 230 [193, 266] 58 [38, 75] 584 [551, 615] 567 [523, 616] 529 [504, 555] 431 [379, 482] 436 [391, 482]
ball-in-cup-catch 968 [965, 973] 984 [982, 986] 983 [981, 985] 981 [979, 984] 982 [981, 984] 981 [978, 983] 982 [980, 984]
cartpole-balance 998 [997, 1000] 999 [998, 1000] 996 [995, 998] 999 [999, 1000] 999 [998, 999] 998 [998, 999] 999 [999, 999]
cartpole-balance-sparse 1000 [1000, 1000] 999 [1000, 1000] 1000 [1000, 1000] 1000 [1000, 1000] 852 [563, 1141] 991 [973, 1008] 967 [904, 1030]
cartpole-swingup 736 [591, 838] 869 [866, 873] 875 [870, 880] 866 [866, 866] 879 [877, 882] 876 [871, 881] 880 [876, 883]
cartpole-swingup-sparse 702 [560, 792] 573 [333, 806] 845 [839, 849] 798 [780, 818] 840 [827, 852] 825 [795, 854] 848 [848, 849]
cheetah-run 917 [915, 920] 699 [655, 744] 914 [911, 917] 877 [849, 905] 863 [822, 904] 920 [918, 922] 821 [642, 913]
finger-spin 666 [577, 763] 335 [99, 596] 986 [986, 988] 937 [917, 956] 988 [987, 989] 849 [758, 939] 891 [810, 972]
finger-turn-easy 906 [883, 927] 912 [774, 983] 979 [975, 983] 953 [931, 974] 957 [923, 992] 935 [903, 968] 953 [925, 980]
finger-turn-hard 864 [812, 900] 470 [199, 727] 947 [916, 977] 950 [910, 974] 957 [920, 993] 915 [859, 972] 951 [925, 977]
fish-swim 813 [808, 819] 86 [64, 120] 659 [615, 706] 792 [773, 810] 618 [523, 713] 823 [799, 846] 826 [806, 846]
hopper-hop 116 [66, 165] 87 [25, 160] 425 [368, 500] 251 [195, 301] 295 [273, 316] 385 [322, 449] 290 [233, 348]
hopper-stand 747 [669, 806] 670 [466, 829] 952 [944, 958] 951 [948, 955] 949 [941, 957] 929 [900, 957] 944 [926, 962]
pendulum-swingup 774 [740, 802] 500 [251, 743] 846 [830, 862] 748 [597, 829] 829 [795, 864] 737 [575, 899] 827 [805, 849]
quadruped-run 130 [92, 169] 645 [567, 713] 942 [938, 947] 947 [940, 954] 859 [824, 895] 928 [916, 939] 935 [928, 943]
quadruped-walk 193 [137, 243] 949 [939, 957] 963 [959, 967] 963 [959, 967] 958 [949, 967] 957 [951, 963] 962 [955, 969]
reacher-easy 966 [964, 970] 970 [951, 982] 983 [980, 986] 983 [983, 985] 983 [983, 984] 983 [981, 986] 983 [979, 986]
reacher-hard 919 [864, 955] 898 [861, 936] 960 [936, 979] 977 [975, 980] 974 [970, 978] 966 [947, 984] 967 [946, 987]
walker-run 510 [430, 588] 804 [783, 825] 854 [851, 859] 793 [765, 815] 790 [776, 805] 796 [792, 801] 817 [812, 821]
walker-stand 941 [934, 948] 983 [974, 989] 991 [990, 994] 988 [987, 990] 990 [986, 994] 985 [982, 989] 987 [984, 990]
walker-walk 898 [875, 919] 977 [975, 980] 981 [979, 984] 978 [978, 980] 979 [975, 983] 975 [972, 978] 976 [974, 978]
IQM 0.813 [0.621, 0.899] 0.771 [0.570, 0.907] 0.941 [0.880, 0.973] 0.927 [0.858, 0.966] 0.928 [0.899, 0.952] 0.922 [0.905, 0.938] 0.933 [0.918, 0.948]
Median 0.813 [0.702, 0.917] 0.804 [0.573, 0.949] 0.952 [0.875, 0.981] 0.950 [0.866, 0.977] 0.872 [0.819, 0.906] 0.870 [0.841, 0.896] 0.875 [0.847, 0.905]
Mean 0.714 [0.584, 0.832] 0.689 [0.548, 0.816] 0.889 [0.819, 0.946] 0.871 [0.788, 0.935] 0.861 [0.823, 0.896] 0.864 [0.84, 0.887] 0.874 [0.849, 0.897]
Figure 16: DMC-Easy Learning Curves. Average episode return for the DMC-Easy environment. Results are averaged over random seeds of each algorithm, and the shaded areas indicate 95% bootstrap confidence intervals.
J.3 Deepmind Control Suite - Hard
Table 18: DMC Hard. Final average performance at 1M environment steps for each of the $7$ tasks of the DMC Hard benchmark. The number of evaluated random seeds for each algorithm is provided in Appendix E. The values in [brackets] represent a 95% bootstrap confidence interval. The aggregate mean, median and interquartile mean (IQM) are reported in units of 1k.
Task DreamerV3 TD7 TD-MPC2 MR.Q Simba SimbaV2
dog-run 4 [4, 5] 69 [36, 101] 265 [166, 342] 569 [547, 595] 544 [525, 564] 562 [516, 608]
dog-stand 22 [20, 27] 582 [432, 741] 506 [266, 715] 967 [960, 975] 960 [951, 969] 981 [977, 985]
dog-trot 10 [6, 17] 21 [13, 30] 407 [265, 530] 877 [845, 898] 824 [773, 876] 861 [772, 950]
dog-walk 17 [15, 21] 52 [19, 116] 486 [240, 704] 916 [908, 924] 916 [905, 928] 935 [927, 944]
humanoid-run 0 [1, 1] 57 [23, 92] 181 [121, 231] 200 [170, 236] 181 [171, 191] 194 [182, 207]
humanoid-stand 5 [5, 6] 317 [117, 516] 658 [506, 745] 868 [822, 903] 846 [801, 890] 916 [886, 945]
humanoid-walk 1 [1, 2] 176 [42, 320] 754 [725, 791] 662 [610, 724] 668 [608, 728] 651 [590, 713]
IQM 0.008 [0.002, 0.016] 0.134 [0.047, 0.343] 0.464 [0.305, 0.632] 0.778 [0.500, 0.911] 0.773 [0.713, 0.83] 0.808 [0.726, 0.879]
Median 0.005 [0.001, 0.018] 0.069 [0.052, 0.317] 0.486 [0.265, 0.658] 0.868 [0.569, 0.916] 0.706 [0.647, 0.772] 0.729 [0.655, 0.808]
Mean 0.009 [0.003, 0.015] 0.182 [0.062, 0.336] 0.465 [0.329, 0.606] 0.723 [0.516, 0.886] 0.706 [0.656, 0.755] 0.729 [0.664, 0.791]
Task PPO SAC SR-SAC iQRL BRO MAD-TD
dog-run 26 [26, 28] 36 [3, 69] 6 [5, 8] 380 [336, 424] 374 [338, 411] 437 [396, 478]
dog-stand 129 [122, 139] 102 [39, 164] 24 [17, 31] 926 [897, 955] 966 [956, 977] 967 [952, 982]
dog-trot 31 [30, 34] 29 [6, 52] 13 [8, 18] 713 [516, 909] 783 [717, 848] 867 [805, 929]
dog-walk 40 [37, 43] 38 [11, 65] 8 [7, 9] 866 [827, 905] 931 [920, 942] 924 [906, 943]
humanoid-run 0 [1, 1] 116 [89, 144] 121 [21, 221] 188 [167, 210] 204 [186, 223] 200 [180, 220]
humanoid-stand 5 [5, 6] 352 [225, 479] 161 [-36, 359] 727 [655, 799] 920 [909, 931] 870 [840, 901]
humanoid-walk 1 [1, 2] 273 [128, 418] 122 [-115, 359] 688 [642, 735] 672 [619, 725] 684 [609, 759]
IQM 0.021 [0.003, 0.069] 0.069 [0.042, 0.114] 0.011 [0.007, 0.036] 0.694 [0.528, 0.805] 0.772 [0.662, 0.854] 0.787 [0.691, 0.865]
Median 0.026 [0.001, 0.040] 0.159 [0.08, 0.183] 0.049 [0.01, 0.095] 0.64 [0.516, 0.766] 0.694 [0.615, 0.774] 0.707 [0.634, 0.786]
Mean 0.033 [0.009, 0.068] 0.136 [0.098, 0.175] 0.065 [0.025, 0.114] 0.642 [0.531, 0.747] 0.693 [0.625, 0.757] 0.708 [0.642, 0.771]
Figure 17: DMC-Hard Learning Curves. Average episode return for the DMC-Hard environment. Results are averaged over random seeds of each algorithm, and the shaded areas indicate 95% bootstrap confidence intervals.
J.4 MyoSuite
Table 19: MyoSuite. Final average performance at 1M environment steps across each of the $10$ continuous control tasks in the MyoSuite benchmark, including both fixed-goal and randomized-goal (hard) settings. The number of evaluated random seeds for each algorithm is provided in Appendix E. The values in [brackets] represent a 95% bootstrap confidence interval. Performance is measured by the average success rate of each task.
Task DreamerV3 TD7 TD-MPC2 Simba SimbaV2
myo-pen-twirl-hard 53.3 [29.8, 76.9] 12.0 [2.4, 21.6] 40.0 [40.0, 40.0] 77.0 [66.4, 87.6] 93.0 [88.8, 97.2]
myo-pen-twirl 96.7 [90.1, 103.2] 100.0 [100.0, 100.0] 70.0 [11.2, 128.8] 80.0 [53.9, 106.1] 100.0 [100.0, 100.0]
myo-key-turn-hard 0.0 [0.0, 0.0] 0.0 [0.0, 0.0] 0.0 [0.0, 0.0] 7.0 [-3.1, 17.1] 62.0 [42.7, 81.3]
myo-key-turn 88.9 [67.1, 110.7] 100.0 [100.0, 100.0] 100.0 [100.0, 100.0] 100.0 [100.0, 100.0] 100.0 [100.0, 100.0]
myo-obj-hold-hard 9.4 [8.4, 10.5] 10.0 [-0.7, 20.7] 56.7 [39.4, 74.0] 96.0 [92.8, 99.2] 98.0 [95.4, 100.6]
myo-obj-hold 33.3 [-32.0, 98.7] 20.0 [-19.2, 59.2] 100.0 [100.0, 100.0] 90.0 [70.4, 109.6] 100.0 [100.0, 100.0]
myo-pose-hard 0.0 [0.0, 0.0] 0.0 [0.0, 0.0] 0.0 [0.0, 0.0] 0.0 [0.0, 0.0] 0.0 [0.0, 0.0]
myo-pose 100.0 [100.0, 100.0] 0.0 [0.0, 0.0] 100.0 [100.0, 100.0] 100.0 [100.0, 100.0] 100.0 [100.0, 100.0]
myo-reach-hard 0.0 [0.0, 0.0] 14.0 [0.7, 27.3] 83.3 [66.0, 100.6] 93.0 [86.4, 99.6] 94.0 [87.3, 100.7]
myo-reach 100.0 [100.0, 100.0] 100.0 [100.0, 100.0] 100.0 [100.0, 100.0] 100.0 [100.0, 100.0] 100.0 [100.0, 100.0]
IQM 46.6 [16.7, 76.9] 22.3 [4.2, 46.2] 77.5 [50.6, 94.4] 95.2 [82.6, 98.8] 99.0 [96.8, 100.0]
Median 47.0 [29.0, 67.0] 34.0 [21.0, 51.0] 65.0 [47.0, 82.0] 77.0 [66.5, 85.5] 84.5 [78.0, 93.0]
Mean 48.2 [31.8, 64.6] 35.6 [23.8, 48.0] 65.0 [50.3, 78.7] 74.3 [66.3, 81.7] 84.7 [78.2, 90.6]
Figure 18: MyoSuite Learning Curves. Average episode success rate (%) for the MyoSuite environment. Results are averaged over random seeds of each algorithm, and the shaded areas indicate 95% bootstrap confidence intervals.
J.5 Humanoid Bench
Table 20: HumanoidBench. Final average performance at 1M environment steps for each of the $14$ locomotion tasks in the HumanoidBench benchmark. The number of evaluated random seeds for each algorithm is provided in Appendix E. The values in [brackets] represent a 95% bootstrap confidence interval. The aggregate mean, median and interquartile mean (IQM) are computed over the success normalized score as described in Appendix F.4.
Task DreamerV3 TD7 TD-MPC2 CQN-AS Simba SimbaV2
h1-pole-v0 41 [28, 54] 441 [320, 563] 744 [609, 879] 102 [94, 111] 716 [667, 765] 791 [785, 797]
h1-slide-v0 11 [7, 15] 39 [26, 53] 334 [304, 364] 50 [47, 54] 277 [252, 303] 487 [404, 571]
h1-stair-v0 7 [2, 12] 52 [31, 74] 378 [108, 648] 47 [39, 54] 269 [153, 385] 493 [467, 518]
h1-balance-hard-v0 11 [7, 15] 79 [51, 107] 31 [5, 56] 46 [44, 49] 75 [71, 80] 143 [128, 157]
h1-balance-simple-v0 9 [6, 12] 69 [58, 80] 42 [14, 70] 63 [57, 70] 337 [193, 482] 723 [651, 795]
h1-sit-hard-v0 15 [-4, 35] 235 [154, 315] 723 [660, 786] 220 [143, 297] 512 [354, 670] 679 [548, 811]
h1-sit-simple-v0 19 [9, 28] 874 [869, 879] 790 [772, 809] 692 [512, 872] 833 [814, 853] 875 [870, 880]
h1-maze-v0 113 [107, 118] 147 [137, 156] 244 [106, 383] 126 [111, 141] 354 [342, 366] 313 [287, 340]
h1-crawl-v0 248 [176, 319] 582 [563, 600] 962 [959, 965] 564 [544, 585] 923 [904, 942] 946 [933, 959]
h1-hurdle-v0 4 [3, 5] 60 [18, 102] 387 [254, 519] 23 [12, 33] 175 [150, 201] 202 [167, 236]
h1-reach-v0 3203 [2824, 3581] 1409 [998, 1821] 2654 [1951, 3357] 1850 [1232, 2469] 3874 [3220, 4527] 3850 [3272, 4427]
h1-run-v0 4 [2, 6] 91 [54, 128] 778 [763, 793] 22 [9, 34] 232 [185, 279] 415 [307, 524]
h1-stand-v0 15 [7, 22] 433 [138, 727] 798 [779, 817] 439 [213, 665] 772 [701, 843] 814 [770, 857]
h1-walk-v0 8 [1, 16] 33 [22, 45] 814 [813, 815] 56 [37, 74] 550 [391, 709] 845 [840, 850]
IQM 0.007 [0.004, 0.012] 0.134 [0.088, 0.245] 0.734 [0.510, 0.936] 0.094 [0.070, 0.161] 0.521 [0.413, 0.633] 0.799 [0.686, 0.908]
Median 0.021 [0.000, 0.047] 0.284 [0.183, 0.392] 0.696 [0.536, 0.881] 0.226 [0.139, 0.307] 0.598 [0.514, 0.692] 0.781 [0.693, 0.865]
Mean 0.022 [0.000, 0.046] 0.289 [0.207, 0.375] 0.710 [0.562, 0.858] 0.221 [0.157, 0.296] 0.606 [0.536, 0.678] 0.776 [0.705, 0.849]
Figure 19: Humanoidbench Learning Curves. Average episode return for the HumanoidBench environment. Results are averaged over random seeds of each algorithm, and the shaded areas indicate 95% bootstrap confidence intervals. The black dotted line indicates the success score of each tasks (Appendix F.4)
Appendix K Complete Ablation Results
This section presents a per-environment analysis of the design variations discussed in Section 5.4. Each table includes raw scores for individual environments, with [bracketed values] indicating 95% bootstrap confidence intervals. The aggregate mean, median, and interquartile mean (IQM) are calculated based on the differences in normalized scores. To illustrate the magnitude of these differences, we use the following highlight scale:
  • $(\geq 0.1)$
  • $[0.05,0.1)$
  • $[0.02,0.05)$
  • $[-0.02,-0.05)$
  • $[-0.05,-0.1)$
  • $(\leq-0.05)$
K.1 Gym - MuJoCo
Table 21: Mujoco (Input Design). Final average performance at 1M environment steps averaged over 3 seeds. The [bracketed values] represent a 95% bootstrap confidence interval. The aggregate mean, median and interquartile mean (IQM) are computed over the TD3-normalized score as described in Appendix F.1.
Task SimbaV2 No L2 Normalize No Shifting $c_{shift}:1$ Resize Projection
Ant-v4 7429 [7209, 7649] 7267 [7065, 7469] 7203 [6765, 7641] 7367 [7302, 7433] 7834 [7374, 8294]
HalfCheetah-v4 12022 [11640, 12404] 5386 [4901, 5870] 5464 [5178, 5750] 11913 [11241, 12584] 12047 [11315, 12778]
Hopper-v4 4053 [3928, 4178] 3764 [3550, 3978] 3676 [3619, 3734] 3703 [3217, 4189] 3788 [3602, 3974]
Humanoid-v4 10545 [10195, 10896] 9820 [8982, 10658] 10655 [10050, 11259] 10680 [10586, 10775] 10530 [10308, 10753]
Walker2d-v4 6938 [6691, 7185] 5560 [4757, 6364] 5750 [5368, 6131] 6192 [5986, 6399] 6985 [6378, 7591]
IQM 1.637 [1.471, 1.792] 1.448 [1.102, 1.715] 1.463 [1.115, 1.746] 1.552 [1.304, 1.809] 1.645 [1.34, 1.919]
Median 1.616 [1.495, 1.746] 1.364 [1.122, 1.618] 1.411 [1.14, 1.67] 1.573 [1.372, 1.749] 1.623 [1.434, 1.813]
Mean 1.617 [1.514, 1.72] 1.37 [1.137, 1.591] 1.406 [1.164, 1.644] 1.558 [1.39, 1.728] 1.623 [1.45, 1.796]
Table 22: Mujoco (Output Design). Final performance at 1M environment steps averaged over 3 seeds. The [bracketed values] represent a 95% bootstrap confidence interval. The aggregate mean, median and interquartile mean (IQM) are computed over the TD3-normalized score as described in Appendix F.1.
Task SimbaV2 MSE Loss No Reward Scaling No Return Bounding Hard Target
Ant-v4 7429 [7209, 7649] 6195 [5459, 6931] 7087 [6935, 7239] 7622 [7486, 7757] 7373 [7342, 7405]
HalfCheetah-v4 12022 [11640, 12404] 12222 [11753, 12691] 12775 [12608, 12941] 12724 [12133, 13315] 11986 [11434, 12538]
Hopper-v4 4053 [3928, 4178] 3507 [$3$, 3682] 2932 [2007, 3858] 4113 [3999, 4228] 3623 [3445, 3800]
Humanoid-v4 10545 [10195, 10896] 7764 [7227, 8302] 8265 [5867, 10664] 10583 [10506, 10660] 9973 [9763, 10184]
Walker2d-v4 6938 [6691, 7185] 5267 [4667, 5866] 5786 [5116, 6456] 6442 [6112, 6772] 7428 [6463, 8393]
IQM 1.637 [1.474, 1.792] 1.334 [1.162, 1.5] 1.405 [1.236, 1.597] 1.612 [1.367, 1.864] 1.624 [1.32, 1.874]
Median 1.616 [1.495, 1.745] 1.341 [1.202, 1.483] 1.426 [1.201, 1.609] 1.616 [1.453, 1.79] 1.593 [1.399, 1.771]
Mean 1.617 [1.516, 1.719] 1.343 [1.225, 1.462] 1.395 [1.247, 1.544] 1.62 [1.471, 1.773] 1.589 [1.414, 1.757]
Table 23: Mujoco (Training Design). Final performance at 1M environment steps averaged over 3 seeds. The [bracketed values] represent a 95% bootstrap confidence interval. The aggregate mean, median and interquartile mean (IQM) are computed over the TD3-normalized score as described in Appendix F.1.
Task SimbaV2 No LR Decay $s_{init}:1$ $s_{scale}:1$ $\alpha_{init}:0.5$ $\alpha_{scale}:1$
Ant-v4 7429 [7209, 7649] 7553 [6914, 8192] 7429 [7237, 7621] 7296 [7146, 7447] 7552 [7323, 7781] 7258 [6959, 7556]
HalfCheetah-v4 12022 [11640, 12404] 12227 [11760, 12694] 12090 [11758, 12422] 11548 [10470, 12626] 12538 [12472, 12604] 11982 [11585, 12378]
Hopper-v4 4053 [3928, 4178] 3635 [3000, 4269] 4046 [3944, 4148] 4008 [3869, 4146] 3568 [3279, 3857] 4023 [3830, 4216]
Humanoid-v4 10545 [10195, 10896] 9907 [8412, 11401] 9819 [8615, 11023] 10636 [10465, 10807] 10828 [10426, 11229] 8624 [6418, 10831]
Walker2d-v4 6938 [6691, 7185] 6661 [6010, 7311] 6583 [5893, 7274] 6770 [6541, 6998] 6328 [5726, 6929] 6744 [6476, 7013]
IQM 1.637 [1.478, 1.789] 1.556 [1.31, 1.81] 1.588 [1.418, 1.749] 1.613 [1.456, 1.767] 1.563 [1.29, 1.848] 1.53 [1.296, 1.746]
Median 1.616 [1.494, 1.743] 1.541 [1.374, 1.745] 1.605 [1.451, 1.709] 1.606 [1.467, 1.721] 1.584 [1.388, 1.78] 1.546 [1.362, 1.681]
Mean 1.617 [1.518, 1.718] 1.562 [1.393, 1.73] 1.571 [1.467, 1.675] 1.594 [1.488, 1.699] 1.583 [1.405, 1.757] 1.52 [1.377, 1.659]
K.2 Deepmind Control Suite - Easy
Table 24: DMC-Easy (Input Design). Final performance at 1M environment steps averaged over 3 seeds. The [bracketed values] represent a 95% bootstrap confidence interval. The aggregate mean, median and interquartile mean are computed over the default reward.
Task SimbaV2 No L2 Normalize No Shifting $c_{shift}:1$ Resize Projection
acrobot-swingup 436 [391, 482] 385 [364, 406] 399 [276, 522] 466 [402, 530] 293 [184, 401]
ball-in-cup-catch 982 [980, 984] 564 [92, 1036] 586 [145, 1026] 983 [979, 987] 983 [979, 986]
cartpole-balance 999 [999, 999] 999 [999, 999] 999 [999, 999] 999 [999, 999] 999 [999, 999]
cartpole-balance-sparse 967 [904, 1030] 1000 [1000, 1000] 992 [978, 1007] 1000 [1000, 1000] 1000 [1000, 1000]
cartpole-swingup 880 [876, 883] 881 [881, 882] 881 [880, 882] 882 [881, 882] 799 [646, 951]
cartpole-swingup-sparse 848 [848, 849] 838 [829, 846] 829 [796, 862] 846 [843, 848] 803 [719, 888]
cheetah-run 920 [918, 922] 499 [426, 572] 519 [477, 560] 919 [914, 924] 917 [914, 920]
finger-spin 891 [810, 972] 699 [485, 913] 620 [404, 836] 855 [674, 1036] 883 [737, 1030]
finger-turn-easy 953 [925, 980] 922 [858, 986] 924 [831, 1017] 922 [874, 970] 925 [866, 985]
finger-turn-hard 951 [925, 977] 888 [796, 980] 937 [895, 979] 870 [861, 879] 917 [864, 970]
fish-swim 826 [806, 846] 450 [316, 584] 442 [326, 558] 806 [783, 830] 758 [706, 811]
hopper-hop 290 [233, 348] 347 [270, 424] 208 [132, 284] 212 [79, 345] 350 [191, 509]
hopper-stand 944 [926, 962] 804 [578, 1031] 680 [444, 915] 925 [883, 967] 753 [440, 1066]
pendulum-swingup 827 [805, 849] 610 [208, 1011] 620 [213, 1026] 810 [762, 859] 678 [453, 902]
quadruped-run 935 [928, 943] 937 [923, 950] 917 [902, 932] 900 [850, 949] 935 [927, 943]
quadruped-walk 962 [955, 969] 962 [949, 975] 963 [958, 967] 960 [945, 974] 957 [947, 967]
reacher-easy 983 [979, 986] 949 [905, 992] 970 [952, 989] 984 [982, 985] 982 [980, 985]
reacher-hard 967 [946, 987] 881 [707, 1054] 935 [890, 979] 973 [966, 979] 975 [972, 978]
walker-run 817 [812, 821] 762 [672, 853] 793 [778, 809] 816 [811, 820] 813 [809, 816]
walker-stand 987 [984, 990] 987 [984, 990] 987 [983, 991] 986 [978, 995] 979 [967, 991]
walker-walk 976 [974, 978] 974 [968, 981] 978 [975, 981] 976 [973, 980] 969 [963, 975]
IQM 0.933 [0.918, 0.948] 0.874 [0.795, 0.922] 0.871 [0.789, 0.921] 0.919 [0.891, 0.942] 0.92 [0.888, 0.946]
Median 0.875 [0.847, 0.905] 0.787 [0.717, 0.839] 0.781 [0.712, 0.837] 0.867 [0.818, 0.905] 0.844 [0.792, 0.892]
Mean 0.874 [0.849, 0.898] 0.779 [0.722, 0.832] 0.771 [0.713, 0.826] 0.862 [0.819, 0.9] 0.842 [0.794, 0.885]
Table 25: DMC-Easy (Output Design). Final performance at 1M environment steps averaged over 3 seeds. The [bracketed values] represent a 95% bootstrap confidence interval. The aggregate mean, median and interquartile mean are computed over the default reward.
Task SimbaV2 MSE Loss No Reward Scaling No Return Bounding Hard Target
acrobot-swingup 436 [391, 482] 384 [273, 494] 383 [329, 438] 439 [386, 492] 423 [350, 496]
ball-in-cup-catch 982 [980, 984] 982 [979, 985] 981 [978, 984] 983 [979, 986] 983 [979, 986]
cartpole-balance 999 [999, 999] 999 [998, 1000] 999 [999, 999] 694 [659, 730] 999 [999, 999]
cartpole-balance-sparse 967 [904, 1030] 1000 [1000, 1000] 970 [913, 1027] 998 [994, 1001] 1000 [1000, 1000]
cartpole-swingup 880 [876, 883] 881 [880, 882] 881 [880, 881] 758 [737, 779] 881 [880, 882]
cartpole-swingup-sparse 848 [848, 849] 845 [843, 847] 715 [493, 937] 844 [839, 849] 846 [844, 848]
cheetah-run 920 [918, 922] 796 [563, 1030] 869 [813, 925] 887 [829, 946] 905 [873, 937]
finger-spin 891 [810, 972] 959 [937, 980] 824 [684, 963] 774 [632, 916] 954 [919, 989]
finger-turn-easy 953 [925, 980] 973 [965, 981] 916 [873, 959] 970 [964, 976] 970 [964, 977]
finger-turn-hard 951 [925, 977] 886 [793, 978] 853 [761, 945] 917 [861, 972] 966 [958, 973]
fish-swim 826 [806, 846] 838 [827, 849] 844 [834, 853] 819 [790, 847] 809 [792, 826]
hopper-hop 290 [233, 348] 380 [198, 561] 294 [237, 350] 211 [114, 309] 316 [293, 339]
hopper-stand 944 [926, 962] 920 [858, 982] 928 [897, 959] 710 [464, 956] 918 [862, 973]
pendulum-swingup 827 [805, 849] 773 [754, 792] 814 [776, 852] 809 [763, 854] 809 [760, 858]
quadruped-run 935 [928, 943] 938 [914, 962] 929 [901, 956] 923 [903, 944] 943 [930, 956]
quadruped-walk 962 [955, 969] 955 [943, 967] 968 [960, 976] 959 [949, 968] 964 [953, 974]
reacher-easy 983 [979, 986] 968 [944, 992] 966 [938, 994] 983 [981, 985] 983 [982, 985]
reacher-hard 967 [946, 987] 978 [977, 979] 976 [972, 981] 976 [972, 981] 969 [949, 989]
walker-run 817 [812, 821] 795 [793, 797] 818 [815, 820] 732 [687, 777] 817 [812, 821]
walker-stand 987 [984, 990] 992 [990, 993] 987 [983, 991] 945 [912, 978] 986 [979, 994]
walker-walk 976 [974, 978] 977 [974, 980] 974 [970, 978] 972 [966, 979] 978 [976, 981]
IQM 0.933 [0.918, 0.948] 0.928 [0.894, 0.954] 0.914 [0.89, 0.937] 0.887 [0.843, 0.922] 0.938 [0.911, 0.96]
Median 0.875 [0.846, 0.905] 0.866 [0.813, 0.918] 0.853 [0.808, 0.896] 0.823 [0.777, 0.871] 0.878 [0.836, 0.918]
Mean 0.874 [0.848, 0.898] 0.868 [0.821, 0.909] 0.852 [0.814, 0.888] 0.824 [0.78, 0.865] 0.878 [0.838, 0.913]
Table 26: DMC-Easy (Training Design). Final performance at 1M environment steps averaged over 3 seeds. The [bracketed values] represent a 95% bootstrap confidence interval. The aggregate mean, median and interquartile mean are computed over the default reward.
Task SimbaV2 No LR Decay $s_{init}:1$ $s_{scale}:1$ $\alpha_{init}:0.5$ $\alpha_{scale}:1$
acrobot-swingup 436 [391, 482] 393 [269, 517] 490 [453, 527] 448 [408, 488] 452 [359, 545] 449 [391, 506]
ball-in-cup-catch 982 [980, 984] 982 [978, 986] 982 [980, 984] 982 [980, 984] 983 [979, 987] 983 [979, 987]
cartpole-balance 999 [999, 999] 999 [999, 999] 999 [999, 999] 999 [999, 999] 999 [999, 999] 999 [999, 999]
cartpole-balance-sparse 967 [904, 1030] 1000 [1000, 1000] 999 [997, 1000] 1000 [1000, 1000] 1000 [1000, 1000] 1000 [1000, 1000]
cartpole-swingup 880 [876, 883] 880 [878, 882] 881 [881, 882] 882 [881, 882] 881 [881, 882] 881 [881, 882]
cartpole-swingup-sparse 848 [848, 849] 699 [409, 989] 847 [846, 848] 787 [697, 878] 848 [847, 849] 847 [845, 849]
cheetah-run 920 [918, 922] 917 [913, 922] 917 [913, 921] 914 [903, 925] 918 [914, 921] 866 [758, 974]
finger-spin 891 [810, 972] 871 [716, 1025] 957 [941, 973] 895 [813, 977] 950 [921, 979] 928 [852, 1004]
finger-turn-easy 953 [925, 980] 895 [846, 943] 945 [904, 986] 955 [931, 979] 864 [794, 934] 878 [800, 956]
finger-turn-hard 951 [925, 977] 869 [755, 982] 941 [910, 971] 959 [937, 981] 939 [895, 983] 943 [898, 988]
fish-swim 826 [806, 846] 805 [766, 844] 819 [804, 834] 821 [792, 850] 806 [759, 853] 825 [814, 835]
hopper-hop 290 [233, 348] 321 [305, 338] 282 [217, 347] 304 [265, 343] 171 [59, 283] 313 [293, 333]
hopper-stand 944 [926, 962] 922 [882, 963] 840 [705, 976] 879 [730, 1027] 929 [882, 976] 622 [240, 1004]
pendulum-swingup 827 [805, 849] 810 [763, 858] 827 [806, 848] 820 [797, 843] 811 [766, 857] 812 [767, 857]
quadruped-run 935 [928, 943] 940 [923, 958] 929 [921, 937] 929 [910, 948] 931 [911, 952] 930 [922, 938]
quadruped-walk 962 [955, 969] 967 [962, 973] 966 [962, 971] 954 [939, 969] 953 [946, 961] 954 [948, 959]
reacher-easy 983 [979, 986] 982 [979, 985] 983 [980, 985] 983 [981, 986] 983 [980, 985] 984 [981, 986]
reacher-hard 967 [946, 987] 977 [971, 982] 949 [920, 977] 958 [933, 983] 973 [968, 979] 970 [958, 982]
walker-run 817 [812, 821] 816 [810, 822] 814 [810, 818] 819 [817, 821] 819 [817, 822] 819 [817, 821]
walker-stand 987 [984, 990] 987 [986, 989] 990 [988, 991] 988 [987, 990] 989 [985, 992] 990 [988, 993]
walker-walk 976 [974, 978] 976 [969, 984] 975 [972, 979] 977 [973, 981] 975 [974, 976] 977 [974, 981]
IQM 0.933 [0.918, 0.948] 0.923 [0.894, 0.947] 0.932 [0.916, 0.947] 0.934 [0.918, 0.949] 0.927 [0.901, 0.949] 0.922 [0.892, 0.947]
Median 0.875 [0.847, 0.904] 0.858 [0.811, 0.902] 0.878 [0.847, 0.906] 0.871 [0.842, 0.902] 0.863 [0.818, 0.911] 0.859 [0.809, 0.902]
Mean 0.874 [0.849, 0.897] 0.858 [0.813, 0.896] 0.873 [0.848, 0.897] 0.87 [0.843, 0.894] 0.866 [0.819, 0.905] 0.856 [0.812, 0.895]
K.3 Deepmind Control Suite - Hard
Table 27: DMC-Hard (Input Design). Final performance at 1M environment steps averaged over 3 seeds. The [bracketed values] represent a 95% bootstrap confidence interval. The aggregate mean, median and interquartile mean are computed over the default reward.
Task SimbaV2 No L2 Normalize No Shifting $c_{shift}:1$ Resize Projection
dog-run 562 [516, 608] 573 [515, 632] 530 [415, 644] 610 [575, 645] 541 [418, 664]
dog-stand 981 [977, 985] 953 [931, 975] 963 [955, 971] 964 [939, 989] 977 [970, 985]
dog-trot 861 [772, 950] 813 [747, 878] 823 [753, 892] 772 [640, 904] 875 [842, 907]
dog-walk 935 [927, 944] 895 [868, 922] 902 [894, 909] 938 [933, 943] 914 [899, 929]
humanoid-run 194 [182, 207] 189 [173, 205] 236 [213, 259] 182 [169, 196] 204 [177, 231]
humanoid-stand 916 [886, 945] 716 [341, 1092] 913 [895, 932] 876 [794, 958] 904 [886, 921]
humanoid-walk 651 [590, 713] 756 [695, 817] 700 [553, 846] 683 [541, 824] 621 [585, 658]
IQM 0.808 [0.726, 0.88] 0.789 [0.65, 0.868] 0.805 [0.667, 0.893] 0.783 [0.663, 0.882] 0.795 [0.659, 0.894]
Median 0.729 [0.655, 0.808] 0.732 [0.595, 0.816] 0.73 [0.619, 0.826] 0.717 [0.606, 0.823] 0.724 [0.61, 0.825]
Mean 0.729 [0.665, 0.79] 0.7 [0.6, 0.795] 0.724 [0.629, 0.814] 0.718 [0.619, 0.809] 0.72 [0.619, 0.811]
Table 28: DMC-Hard (Output Design). Final performance at 1M environment steps averaged over 3 seeds. The [bracketed values] represent a 95% bootstrap confidence interval. The aggregate mean, median and interquartile mean are computed over the default reward.
Task SimbaV2 MSE Loss No Reward Scaling No Return Bounding Hard Target
dog-run 562 [516, 608] 545 [450, 639] 478 [402, 554] 617 [537, 696] 676 [654, 699]
dog-stand 981 [977, 985] 976 [959, 993] 967 [956, 978] 969 [958, 981] 980 [971, 989]
dog-trot 861 [772, 950] 841 [796, 886] 737 [614, 859] 884 [803, 964] 848 [763, 934]
dog-walk 935 [927, 944] 905 [883, 928] 925 [915, 935] 922 [899, 945] 928 [891, 964]
humanoid-run 194 [182, 207] 173 [146, 200] 237 [181, 293] 182 [154, 209] 209 [159, 260]
humanoid-stand 916 [886, 945] 786 [612, 960] 879 [821, 936] 851 [744, 958] 928 [920, 937]
humanoid-walk 651 [590, 713] 729 [577, 880] 754 [643, 865] 706 [563, 849] 645 [602, 689]
IQM 0.808 [0.728, 0.881] 0.78 [0.62, 0.877] 0.778 [0.655, 0.871] 0.812 [0.69, 0.901] 0.814 [0.706, 0.902]
Median 0.729 [0.655, 0.809] 0.705 [0.57, 0.836] 0.715 [0.61, 0.817] 0.731 [0.621, 0.84] 0.746 [0.641, 0.847]
Mean 0.729 [0.665, 0.79] 0.708 [0.586, 0.813] 0.712 [0.626, 0.792] 0.733 [0.632, 0.824] 0.746 [0.648, 0.833]
Table 29: DMC-Hard (Training Design). Final performance at 1M environment steps averaged over 3 seeds. The [bracketed values] represent a 95% bootstrap confidence interval. The aggregate mean, median and interquartile mean are computed over the default reward.
Task SimbaV2 No LR Decay $s_{init}:1$ $s_{scale}:1$ $\alpha_{init}:0.5$ $\alpha_{scale}:1$
dog-run 562 [516, 608] 516 [407, 625] 586 [550, 621] 543 [488, 598] 556 [420, 691] 546 [493, 599]
dog-stand 981 [977, 985] 957 [933, 980] 976 [970, 982] 972 [964, 979] 951 [912, 989] 979 [969, 989]
dog-trot 861 [772, 950] 770 [663, 878] 836 [756, 916] 850 [798, 901] 870 [842, 898] 814 [692, 937]
dog-walk 935 [927, 944] 927 [919, 935] 938 [921, 955] 949 [937, 960] 924 [903, 945] 931 [921, 942]
humanoid-run 194 [182, 207] 221 [165, 277] 187 [178, 196] 194 [177, 211] 193 [170, 217] 182 [172, 192]
humanoid-stand 916 [886, 945] 932 [916, 948] 874 [812, 936] 823 [749, 897] 900 [877, 924] 918 [892, 944]
humanoid-walk 651 [590, 713] 706 [590, 822] 624 [592, 656] 610 [589, 630] 697 [583, 812] 622 [611, 633]
IQM 0.808 [0.725, 0.879] 0.798 [0.658, 0.895] 0.783 [0.708, 0.851] 0.766 [0.687, 0.838] 0.819 [0.685, 0.892] 0.781 [0.642, 0.891]
Median 0.729 [0.655, 0.808] 0.716 [0.616, 0.822] 0.719 [0.646, 0.794] 0.711 [0.634, 0.782] 0.724 [0.62, 0.833] 0.715 [0.603, 0.823]
Mean 0.729 [0.664, 0.791] 0.719 [0.623, 0.809] 0.718 [0.656, 0.777] 0.706 [0.644, 0.767] 0.728 [0.627, 0.819] 0.714 [0.61, 0.81]
K.4 Myosuite
Table 30: Myosuite (Input Design). Final performance at 1M environment steps averaged over 3 seeds. The [bracketed values] represent a 95% bootstrap confidence interval. The aggregate mean, median and interquartile mean are computed over the default reward.
Task SimbaV2 No L2 Normalize No Shifting $c_{shift}:1$ Resize Projection
myo-key-turn 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0]
myo-key-turn-hard 0.62 [0.427, 0.813] 0.325 [-0.009, 0.659] 0.25 [-0.044, 0.544] 0.8 [0.661, 0.939] 0.85 [0.752, 0.948]
myo-obj-hold 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0]
myo-obj-hold-hard 0.98 [0.954, 1.006] 0.975 [0.926, 1.024] 0.975 [0.926, 1.024] 1.0 [1.0, 1.0] 0.975 [0.926, 1.024]
myo-pen-twirl 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0]
myo-pen-twirl-hard 0.93 [0.888, 0.972] 0.9 [0.82, 0.98] 0.875 [0.689, 1.061] 0.975 [0.926, 1.024] 0.8 [0.604, 0.996]
myo-pose 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0]
myo-pose-hard 0.0 [0.0, 0.0] 0.0 [0.0, 0.0] 0.0 [0.0, 0.0] 0.0 [0.0, 0.0] 0.0 [0.0, 0.0]
myo-reach 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0]
myo-reach-hard 0.94 [0.873, 1.007] 0.95 [0.893, 1.007] 0.9 [0.82, 0.98] 0.925 [0.831, 1.019] 0.9 [0.82, 0.98]
IQM 0.99 [0.968, 1.0] 0.98 [0.885, 1.0] 0.98 [0.86, 1.0] 1.0 [0.955, 1.0] 0.975 [0.925, 1.0]
Median 0.845 [0.78, 0.925] 0.815 [0.695, 0.935] 0.805 [0.68, 0.92] 0.875 [0.765, 0.975] 0.86 [0.75, 0.955]
Mean 0.847 [0.782, 0.906] 0.815 [0.702, 0.915] 0.8 [0.682, 0.905] 0.87 [0.77, 0.953] 0.852 [0.75, 0.938]
Table 31: Myosuite (Output Design). Final performance at 1M environment steps averaged over 3 seeds. The [bracketed values] represent a 95% bootstrap confidence interval. The aggregate mean, median and interquartile mean are computed over the default reward.
Task SimbaV2 MSE Loss No Reward Scaling No Return Bounding Hard Target
myo-key-turn 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0]
myo-key-turn-hard 0.62 [0.427, 0.813] 0.2 [-0.192, 0.592] 0.76 [0.66, 0.86] 0.225 [-0.153, 0.603] 0.675 [0.417, 0.933]
myo-obj-hold 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0]
myo-obj-hold-hard 0.98 [0.954, 1.006] 0.933 [0.868, 0.999] 0.98 [0.941, 1.019] 1.0 [1.0, 1.0] 0.975 [0.926, 1.024]
myo-pen-twirl 1.0 [1.0, 1.0] 0.667 [0.013, 1.32] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0]
myo-pen-twirl-hard 0.93 [0.888, 0.972] 0.867 [0.605, 1.128] 0.98 [0.941, 1.019] 0.9 [0.82, 0.98] 0.9 [0.82, 0.98]
myo-pose 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 0.8 [0.408, 1.192] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0]
myo-pose-hard 0.0 [0.0, 0.0] 0.0 [0.0, 0.0] 0.0 [0.0, 0.0] 0.0 [0.0, 0.0] 0.0 [0.0, 0.0]
myo-reach 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0]
myo-reach-hard 0.94 [0.873, 1.007] 0.9 [0.787, 1.013] 0.88 [0.766, 0.994] 0.925 [0.831, 1.019] 0.925 [0.831, 1.019]
IQM 0.99 [0.968, 1.0] 0.944 [0.712, 1.0] 0.985 [0.931, 1.0] 0.985 [0.87, 1.0] 0.98 [0.93, 1.0]
Median 0.845 [0.78, 0.93] 0.77 [0.58, 0.94] 0.85 [0.74, 0.96] 0.79 [0.68, 0.93] 0.845 [0.74, 0.955]
Mean 0.847 [0.783, 0.906] 0.757 [0.613, 0.887] 0.84 [0.746, 0.922] 0.805 [0.685, 0.912] 0.848 [0.745, 0.938]
Table 32: Myosuite (Training Design). Final performance at 1M environment steps averaged over 3 seeds. The [bracketed values] represent a 95% bootstrap confidence interval. The aggregate mean, median and interquartile mean are computed over the default reward.
Task SimbaV2 No LR Decay $s_{init}:1$ $s_{scale}:1$ $\alpha_{init}:0.5$ $\alpha_{scale}:1$
myo-key-turn 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 0.9 [0.704, 1.096] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0]
myo-key-turn-hard 0.62 [0.427, 0.813] 0.65 [0.345, 0.955] 0.74 [0.585, 0.895] 0.69 [0.502, 0.878] 0.675 [0.581, 0.769] 0.85 [0.662, 1.038]
myo-obj-hold 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0]
myo-obj-hold-hard 0.98 [0.954, 1.006] 1.0 [1.0, 1.0] 0.95 [0.897, 1.003] 0.98 [0.954, 1.006] 0.95 [0.893, 1.007] 1.0 [1.0, 1.0]
myo-pen-twirl 1.0 [1.0, 1.0] 0.75 [0.26, 1.24] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0]
myo-pen-twirl-hard 0.93 [0.888, 0.972] 0.775 [0.451, 1.099] 0.89 [0.81, 0.97] 0.88 [0.816, 0.944] 0.925 [0.831, 1.019] 1.0 [1.0, 1.0]
myo-pose 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0]
myo-pose-hard 0.0 [0.0, 0.0] 0.0 [0.0, 0.0] 0.0 [0.0, 0.0] 0.0 [0.0, 0.0] 0.0 [0.0, 0.0] 0.0 [0.0, 0.0]
myo-reach 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0] 1.0 [1.0, 1.0]
myo-reach-hard 0.94 [0.873, 1.007] 0.925 [0.831, 1.019] 0.97 [0.928, 1.012] 0.91 [0.848, 0.972] 0.875 [0.752, 0.998] 0.9 [0.82, 0.98]
IQM 0.99 [0.968, 1.0] 0.985 [0.875, 1.0] 0.99 [0.964, 1.0] 0.982 [0.948, 1.0] 0.975 [0.905, 1.0] 1.0 [0.97, 1.0]
Median 0.845 [0.78, 0.925] 0.84 [0.7, 0.94] 0.865 [0.785, 0.935] 0.84 [0.77, 0.92] 0.85 [0.735, 0.945] 0.875 [0.77, 0.98]
Mean 0.847 [0.782, 0.906] 0.81 [0.698, 0.908] 0.855 [0.792, 0.913] 0.836 [0.77, 0.896] 0.842 [0.742, 0.928] 0.875 [0.772, 0.96]
K.5 Humanoid Bench
Table 33: HumanoidBench (Input Design). Final performance at 1M environment steps averaged over 3 seeds. The [bracketed values] represent a 95% bootstrap confidence interval. The aggregate mean, median and interquartile mean are computed over the default reward.
Task SimbaV2 No L2 Normalize No Shifting $c_{shift}:1$ Resize Projection
h1-balance-hard-v0 143 [128, 157] 113 [62, 164] 94 [80, 109] 89 [74, 104] 150 [105, 194]
h1-balance-simple-v0 723 [651, 795] 760 [666, 854] 512 [236, 789] 799 [744, 854] 760 [658, 861]
h1-crawl-v0 946 [933, 959] 901 [870, 932] 838 [813, 864] 948 [927, 968] 884 [723, 1044]
h1-hurdle-v0 202 [167, 236] 177 [168, 187] 172 [162, 183] 191 [160, 222] 218 [152, 285]
h1-maze-v0 313 [287, 340] 337 [313, 361] 302 [230, 374] 341 [337, 345] 338 [311, 365]
h1-pole-v0 791 [785, 797] 746 [722, 769] 759 [754, 764] 784 [781, 786] 805 [795, 816]
h1-reach-v0 3850 [3272, 4427] 4564 [3107, 6020] 3263 [2743, 3783] 5415 [4354, 6475] 3722 [2105, 5339]
h1-run-v0 415 [307, 524] 212 [208, 217] 218 [196, 240] 365 [115, 614] 585 [377, 792]
h1-sit-hard-v0 679 [548, 811] 493 [227, 758] 721 [499, 944] 757 [667, 847] 667 [402, 933]
h1-sit-simple-v0 875 [870, 880] 829 [783, 874] 803 [745, 860] 869 [858, 880] 890 [872, 909]
h1-slide-v0 487 [404, 571] 518 [499, 537] 497 [483, 511] 520 [467, 573] 435 [318, 551]
h1-stair-v0 493 [467, 518] 477 [460, 494] 475 [433, 517] 540 [480, 600] 515 [504, 525]
h1-stand-v0 814 [770, 857] 821 [786, 856] 835 [833, 837] 848 [845, 851] 764 [661, 867]
h1-walk-v0 845 [840, 850] 675 [560, 789] 832 [814, 850] 842 [830, 853] 845 [835, 855]
IQM 0.799 [0.684, 0.907] 0.709 [0.525, 0.881] 0.708 [0.509, 0.908] 0.833 [0.635, 0.997] 0.808 [0.61, 0.977]
Median 0.781 [0.69, 0.862] 0.698 [0.566, 0.851] 0.698 [0.552, 0.85] 0.801 [0.639, 0.944] 0.801 [0.631, 0.934]
Mean 0.776 [0.704, 0.847] 0.711 [0.589, 0.833] 0.7 [0.578, 0.825] 0.791 [0.66, 0.921] 0.779 [0.653, 0.905]
Table 34: HumanoidBench (Output Design). Final performance at 1M environment steps averaged over 3 seeds. The [bracketed values] represent a 95% bootstrap confidence interval. The aggregate mean, median and interquartile mean are computed over the default reward.
Task SimbaV2 MSE Loss No Reward Scaling No Return Bounding Hard Target
h1-balance-hard-v0 143 [128, 157] 89 [73, 106] 92 [82, 101] 136 [117, 155] 173 [133, 213]
h1-balance-simple-v0 723 [651, 795] 765 [664, 865] 779 [680, 877] 832 [825, 839] 739 [553, 924]
h1-crawl-v0 946 [933, 959] 945 [921, 969] 953 [934, 971] 942 [896, 989] 938 [913, 964]
h1-hurdle-v0 202 [167, 236] 243 [241, 245] 140 [105, 174] 192 [132, 252] 207 [186, 228]
h1-maze-v0 313 [287, 340] 324 [282, 367] 273 [220, 327] 320 [291, 350] 339 [309, 370]
h1-pole-v0 791 [785, 797] 774 [773, 776] 724 [625, 823] 796 [791, 800] 787 [786, 788]
h1-reach-v0 3850 [3272, 4427] 5044 [2733, 7355] 4032 [2583, 5480] 4396 [2059, 6732] 3984 [2231, 5738]
h1-run-v0 415 [307, 524] 485 [199, 771] 359 [150, 568] 360 [234, 487] 341 [175, 508]
h1-sit-hard-v0 679 [548, 811] 611 [350, 873] 729 [648, 811] 727 [681, 773] 684 [362, 1005]
h1-sit-simple-v0 875 [870, 880] 869 [864, 875] 871 [859, 883] 868 [863, 872] 871 [870, 871]
h1-slide-v0 487 [404, 571] 352 [260, 445] 410 [313, 506] 534 [484, 583] 581 [550, 613]
h1-stair-v0 493 [467, 518] 512 [470, 554] 450 [382, 517] 510 [495, 526] 515 [495, 536]
h1-stand-v0 814 [770, 857] 742 [621, 862] 775 [684, 865] 751 [570, 933] 645 [251, 1038]
h1-walk-v0 845 [840, 850] 851 [840, 863] 742 [556, 928] 845 [834, 855] 838 [835, 842]
IQM 0.799 [0.684, 0.907] 0.778 [0.594, 0.95] 0.745 [0.584, 0.896] 0.819 [0.623, 0.977] 0.778 [0.572, 0.965]
Median 0.781 [0.692, 0.863] 0.762 [0.614, 0.917] 0.722 [0.605, 0.864] 0.776 [0.636, 0.933] 0.772 [0.614, 0.923]
Mean 0.776 [0.704, 0.848] 0.767 [0.637, 0.894] 0.735 [0.629, 0.842] 0.787 [0.658, 0.915] 0.77 [0.64, 0.897]
Table 35: HumanoidBench (Training Design). Final performance at 1M env steps averaged over 3 seeds. The [bracketed values] represent a 95% bootstrap confidence interval. The aggregate mean, median and interquartile mean are computed over the default reward.
Task SimbaV2 No LR Decay $s_{init}:1$ $s_{scale}:1$ $\alpha_{init}:0.5$ $\alpha_{scale}:1$
h1-balance-hard-v0 143 [128, 157] 118 [86, 150] 139 [119, 159] 139 [116, 162] 152 [140, 164] 139 [119, 159]
h1-balance-simple-v0 723 [651, 795] 812 [758, 867] 815 [794, 836] 813 [793, 833] 763 [620, 907] 732 [645, 820]
h1-crawl-v0 946 [933, 959] 955 [933, 977] 956 [947, 965] 946 [929, 963] 933 [911, 954] 962 [954, 970]
h1-hurdle-v0 202 [167, 236] 150 [64, 235] 228 [218, 238] 188 [159, 218] 203 [187, 219] 209 [204, 214]
h1-maze-v0 313 [287, 340] 283 [150, 417] 347 [339, 355] 337 [318, 356] 367 [352, 382] 333 [324, 342]
h1-pole-v0 791 [785, 797] 736 [625, 847] 774 [760, 788] 767 [736, 798] 791 [777, 804] 792 [785, 798]
h1-reach-v0 3850 [3272, 4427] 5986 [3542, 8430] 4295 [3423, 5167] 4988 [4251, 5724] 4312 [2719, 5905] 5495 [3782, 7207]
h1-run-v0 415 [307, 524] 424 [172, 675] 357 [277, 436] 351 [246, 457] 234 [209, 259] 337 [171, 503]
h1-sit-hard-v0 679 [548, 811] 678 [382, 975] 686 [575, 796] 734 [604, 863] 507 [257, 757] 702 [493, 911]
h1-sit-simple-v0 875 [870, 880] 845 [793, 897] 875 [869, 881] 877 [871, 882] 868 [852, 885] 870 [868, 872]
h1-slide-v0 487 [404, 571] 388 [268, 507] 486 [420, 553] 525 [475, 574] 447 [380, 514] 533 [512, 554]
h1-stair-v0 493 [467, 518] 512 [499, 525] 507 [491, 523] 504 [483, 525] 513 [505, 521] 503 [468, 538]
h1-stand-v0 814 [770, 857] 603 [434, 773] 751 [697, 804] 789 [700, 879] 793 [707, 878] 799 [720, 877]
h1-walk-v0 845 [840, 850] 843 [832, 855] 843 [837, 849] 839 [828, 850] 834 [829, 839] 827 [802, 851]
IQM 0.799 [0.683, 0.908] 0.764 [0.578, 0.93] 0.796 [0.688, 0.901] 0.823 [0.709, 0.927] 0.725 [0.532, 0.917] 0.817 [0.615, 0.978]
Median 0.781 [0.69, 0.862] 0.769 [0.599, 0.912] 0.776 [0.7, 0.866] 0.792 [0.705, 0.876] 0.746 [0.595, 0.893] 0.802 [0.64, 0.938]
Mean 0.776 [0.704, 0.847] 0.754 [0.625, 0.883] 0.781 [0.711, 0.852] 0.789 [0.719, 0.861] 0.745 [0.619, 0.872] 0.792 [0.659, 0.916]